### v1

# EngChain: A Symbolic Benchmark for Verifiable Multi-Step Reasoning in Engineering

Abstract
Large Language Models (LLMs) are increasingly being applied to specialized, high-stakes domains like engineering, which demands rigorous evaluation of their complex reasoning capabilities. While current benchmarks assess language understanding, factual recall, mathematics or code generation, none capture the integrative reasoning central to engineering where scientific principles, quantitative modeling and practical constraints must converge. To address this gap, we introduce EngChain, a benchmark for verifiable multi-step engineering problem-solving. EngChain contains 90 problems spanning three engineering branches, organized into 9 domains and 20 distinct areas. The problems are generated from symbolic templates with a high degree of randomization to ensure diversity and eliminate the risk of contamination. With this benchmark, we move beyond final answer accuracy with a two-stage evaluation: we first quantitatively verify the numerical and semantic validity of each reasoning step and then introduce LLM-as-a-Judge, an automated system to qualitatively categorize the identified reasoning errors.
EngChain: A Symbolic Benchmark for Verifiable Multi-Step Reasoning in Engineering
Ayesha Gull1 Muhammad Usman Safder1 Rania Elbadry2 Preslav Nakov2 Zhuohan Xie2 1Namal University 2MBZUAI ProjectCodeApp
**footnotetext: Equal contribution.
1 Introduction
As large language models (LLMs) transition from general-purpose language tasks to specialized, high-stakes domains such as scientific discovery, industrial design, and engineering analysis, the need for rigorous and meaningful evaluation has become significant (Zhao et al., 2023; Xie et al., 2023a). The progress of the field is gated by the quality of its benchmarks: without tests that capture practical, physically grounded reasoning, it is difficult to separate genuine capability from sophisticated mimicry.
To meet this challenge, we introduce EngChain, a new benchmark for verifiable, multi-step engineering problem-solving. Our benchmark is built on a foundation of symbolic templates, each capable of programmatically generating a vast number of unique, verifiable problems. Figure 1 provides a concrete example from our benchmark, illustrating how a single template generates a complete problem instance along with its full chain-of-thought solution.
Figure 1: An example from EngChain (CSTR Volume Calculation). A symbolic template (1) generates a unique problem instance (2) and its verifiable, step-by-step chain-of-thought solution (3).
This design directly targets two key problems in current LLM evaluation. The first is “benchmark saturation” (Ott et al., 2022), where models rapidly achieve superhuman performance on static benchmarks like GLUE (Wang et al., 2019) and even their more complex successors such as BBH (Kazemi et al., 2025). The second is a structural limitation: most benchmarks evaluate critical skills in disciplinary silos. General knowledge benchmarks (e.g., MMLU (Hendrycks et al., 2020)) test broad factual recall, while specialized benchmarks (e.g., MATH (Hendrycks et al., 2021b) or HumanEval (Chen et al., 2021)) assess abstract logic and algorithmic translation, respectively.
This fragmentation is limiting for engineering which is fundamentally an integrative discipline that requires the synthesis of these isolated skills, combining scientific principles, mathematical modeling, and practical constraints (e.g., Moaveni, 2019; Dym et al., 2012). This holistic, procedural reasoning is not captured by existing benchmarks (e.g., Felten et al., 2025; Li et al., 2025b), which predominantly focus on final-answer accuracy. In engineering, where a flawed process can lead to catastrophic failure, this evaluation gap is critical. To address this gap, EngChain uses symbolic templates to produce integrative reasoning problems and a two-stage evaluation via LLM-as-a-Judge that verifies the full reasoning chain rather than just the final answer.
Verifiable multi-step reasoning is fundamental to building trustworthy AI systems across all domains. Concurrent with our work, FinChain (Xie et al., 2025) introduced a symbolic benchmark for verifiable chain-of-thought financial reasoning. However, the two benchmarks address fundamentally different reasoning types. FinChain evaluates reasoning based on economic principles, whereas EngChain evaluates reasoning grounded in the physical laws and practical constraints of engineering. Moreover, FinChain uses a global alignment metric to score the entire reasoning path, while EngChain employs a two-stage evaluation that provides both a quantitative Step F1-Score for procedural correctness and a qualitative, diagnostic error analysis using our LLM-as-a-Judge system.
In summary, our contributions are threefold:
  1. 1.
    We introduce EngChain, a new symbolic benchmark for verifiable, multi-step engineering problem-solving with 90 problem templates across nine domains.
  2. 2.
    We propose a rigorous two-stage evaluation framework, including our qualitative LLM-as-a-Judge system, that moves beyond final answer accuracy to diagnose reasoning failures.
  3. 3.
    We present a comprehensive analysis of 11 frontier LLMs, revealing a critical “right for the wrong reasons" gap and identifying that the majority of reasoning failures are due to Conceptual Errors, not calculation slips.
2 Related Work
2.1 Benchmarks in Mathematics and Coding
A large body of work on multi-step reasoning benchmarks has focused on foundational, yet isolated, cognitive skills, particularly in mathematics and coding. In mathematical reasoning, benchmarks ranging from grade-school problems (GSM8K)  (Cobbe et al., 2021) to high-school and college-level competition problems (MATH)  (Hendrycks et al., 2021b) establish a clear standard for evaluating abstract logical deduction. This is further refined by HARDMATH  (Fan et al., 2024) for applied mathematics and GSM-Symbolic  (Mirzadeh et al., 2024) for programming and symbolic tasks. Similarly, benchmarks for coding and software engineering evaluate models across different levels of complexity. HumanEval  (Chen et al., 2021) focuses on translating natural language instructions into small, correct code snippets. MBPP  (Austin et al., 2021) and its derivative MBUPP extend this by testing models on a broader set of programming problems with varying difficulty. SWE-Bench  (Jimenez et al., 2024) addresses complex, real-world codebases, measuring a model’s ability to solve practical software engineering challenges. While these benchmarks are essential for capturing foundational skills, their focus on isolated abilities limits their usefulness for engineering. A model can excel at solving abstract equations or writing pure algorithms yet still be unable to apply those skills to a problem constrained by physical laws and material properties.
2.2 Benchmarks in the Physical Sciences
Recognizing this limitation, later benchmarks focus on applied reasoning within the physical sciences. UGPhysics  (Xu et al., 2025) provides broad coverage and tests how models apply mathematical reasoning to physical laws. Building on this, PhysReason  (Zhang et al., 2025a) emphasizes process-oriented problem solving, guiding models through step-by-step solutions. ABench-Physics  (Zhang et al., 2025b) introduces highly challenging tasks to evaluate advanced capabilities, while PHYBench  (Qiu et al., 2025) targets symbolic computation. NEWTON  (Wang et al., 2023) assesses qualitative understanding, complementing the more quantitative approaches. Other benchmarks, such as LLM-SRBench  (Shojaee et al., 2025), focus specifically on quantitative scientific modeling. Together, these benchmarks show how models apply mathematical reasoning to well-defined physical laws, marking an important advance towards assessing real-world problem-solving. However, they remain largely confined to physics and have not been widely adopted for evaluating models in real-world engineering practices.
2.3 Benchmarks in Engineering
Current approaches to evaluating engineering reasoning are fractured, split between benchmarks that are overly broad and those that are deep but methodologically limited. On one hand, generalist benchmarks like MMLU  (Hendrycks et al., 2021a) and BIG-Bench  (Luo et al., 2024) include engineering topics, but their reliance on multiple-choice, knowledge-based questions primarily tests factual recall rather than the deep, procedural reasoning required for practical engineering problem-solving. On the other hand, more specialized engineering benchmarks, while deeper in scope, either remain confined to a single scientific discipline or lack a scalable, deterministic method for verifying the reasoning process itself. For example, EngiBench  (Felten et al., 2025) provides broad coverage but uses rubric-based scoring for its more complex problems, whereas FEA-Bench  (Li et al., 2025b) evaluates a model’s ability to use professional simulation software, yet this measures tool proficiency rather than fundamental reasoning.
The preceding review makes it clear that the field lacks a benchmark that captures the integrative, cross-disciplinary, and verifiable reasoning central to engineering. To address this critical gap, we introduce EngChain, a benchmark for verifiable, multi-step engineering reasoning. It uses a two-stage evaluation that combines quantitative verification of each intermediate step with a qualitative error diagnosis, covering 90 distinct problem types across nine domains and three engineering branches.
3 EngChain Taxonomy
EngChain covers a wide range of core engineering problems, focusing on domains that are central to real-world practice, which we show in Figure 2.
In Chemical Engineering, it targets the three pillars of chemical process design and analysis. Kinetics problems test the modeling of reaction rates and mechanisms, requiring reasoning about how systems evolve over time. Thermodynamics problems center on energy balances and phase behavior, capturing the constraints that govern feasibility and efficiency. Transport phenomena problems address momentum, heat, and mass transfer, linking microscopic principles to large-scale process behavior. In Electrical Engineering, it covers the full spectrum from fundamental physics to applied systems. At the foundation, it includes electromagnetics, which governs the behavior of fields and waves. Building on this, it addresses signals and systems, focusing on representation, transformation, and filtering of information. Finally, it extends to communications, where these principles are applied to the reliable transmission and processing of signals in real-world contexts. In Mechanical Engineering, it spans three core domains. Fluid mechanics problems address laminar and turbulent flow, testing how models capture momentum and energy transfer in fluids. Mechanics of materials emphasizes stress–strain relationships, deformation, and failure analysis, reflecting how structures respond to applied loads. Vibrations problems focus on natural frequencies and damping, capturing the dynamic stability of mechanical systems. Together, these areas test how forces, motion, and stability interact in physical systems.
Figure 2: The taxonomy of EngChain.
By integrating these principle subfields across the main branches of engineering, EngChain provides a comprehensive foundation of the analytical and procedural reasoning skills essential for engineering problem-solving.
4 EngChain Design and Methodology
To ensure the rigor, scalability, and long-term viability of the benchmark, we adopted a programmatic, template-based methodology. Unlike static datasets, which are finite and susceptible to being inadvertently included in the training corpora of future models, a programmatic approach offers several key advantages. It provides virtually limitless scalability and diversity, as a single template can generate thousands of unique problem instances. This design also ensures strong resistance to training data contamination, as it tests for generalizable reasoning on novel problems rather than the memorization of previously seen solutions.
4.1 Content Selection Framework
The selection of content for EngChain followed a principled, multi-stage methodology designed to ensure the benchmark is both representative of a formal engineering curriculum and grounded in foundational, real-world principles. The process involved a hierarchical selection of domains, areas, and finally, the specific problems to be templatized.
4.1.1 Domain and Area Selection
The nine core domains that form the backbone of EngChain were identified using a two-pronged approach. First, a thorough literature review was conducted, surveying established engineering taxonomies and curricular standards set by leading professional and accreditation bodies. This involved analyzing the program criteria from organizations like the Accreditation Board for Engineering and Technology (ABET, 2024), which mandates core competencies in areas like thermodynamics and circuit analysis for accredited degrees. We also reviewed the educational frameworks of major professional societies, such as the American Society of Mechanical Engineers (ASME, 2025), the Institute of Electrical and Electronics Engineers (IEEE, 2025), and the American Institute of Chemical Engineers (AIChE, 2025), to ensure our selected domains represent a consensus on the foundational knowledge required for professional practice.
Second, these findings were cross-referenced by querying an LLM (gemini-2.0-flash). The model was prompted to identify and rank the most fundamental, cornerstone domains within each of the three primary engineering branches. The full prompt used for this domain-level validation is provided in Appendix A. The final set of nine domains represents a strong consensus between traditional academic structures and the knowledge synthesized by modern AI. A similar two-pronged methodology was then used to identify the most critical areas within each domain: first, by extracting the core pedagogical units from the authoritative textbooks, and second, by cross-referencing this list using a LLM (gemini-2.0-flash). The full prompt used for this area-level validation is provided in Appendix B.
4.1.2 Template Selection
Once the domains and areas were defined, we developed the 90 problem templates. The core engineering principles, problem typologies, and typical constraints for these templates were sourced from the authoritative and widely-adopted textbooks in each field. This textbook-driven approach ensures that every problem in EngChain is pedagogically sound and central to engineering practice. A complete list of these foundational texts is provided in Appendix C.
Crucially, the number of templates developed for each area was not uniform. Instead, it was guided by a Pedagogical Significance Score assigned to each area. This score was programmatically determined by querying an expert-persona LLM (gemini-2.0-flash) to rate each area’s centrality to the curriculum on a 1-5 scale, as detailed in Appendix D. Foundational concepts with higher scores received more extensive coverage. For example, within the Reaction Kinetics domain, the area of Mole Balances (Significance Score: 5) was covered by five templates, while the more specialized application of Levenspiel Plot Interpretation (Significance Score: 1) was covered by a single template. Therefore, the varying number of templates per area is a direct and intentional outcome of our methodology. This ensures that EngChain’s structure reflects the hierarchical and foundational nature of engineering knowledge as it is taught and practiced.
4.2 Template Generation Pipeline
The template generation process follows a systematic pipeline as shown in Figure 3. First, we generate domain-aware parameters are generated. Next, the core engineering computations are performed. These results are then thoroughly validated to ensure they are physically plausible. Finally, the validated content is assembled into a natural language question and its corresponding step-by-step solution.
Figure 3: Our template generation pipeline.
We format the output strings to be both human-readable, like a textbook example, and machine-parsable for automated evaluation. After template functions are defined, we generate the problem instances by repeatedly calling each template, with every run using a unique random seed for reproducibility.
4.3 Template Structure and Outputs
The core of our benchmark is the template function, a standardized Python function that serves as the blueprint for a specific type of engineering problem. To ensure consistency, every template sticks to a simple, standardized signature i.e., it takes no arguments and returns a (question, solution) tuple. The question is a formatted string containing the problem statement to be presented to the LLM, whereas the solution is a formatted string containing the complete reasoning trace rather than only a numerical answer. A generic multi-disciplinary template structure is provided in Figure 4.
Figure 4: Generic multi-disciplinary template structure.
Each solution is structured like a textbook example such that the principles are stated, values are substituted, and intermediate steps are shown with consistent units.
4.4 Domain-Aware Parameterization and Sanity Checks
EngChain uses domain-aware parameterization, ensuring that problems are not just mathematically solvable but are also grounded in physical and engineering realism. This is done by constraining all generated values by the principles of the domain being tested.
For instance, in Chemical Engineering, problems use real reactants like Benzene, pre-validated stoichiometric equations, and temperature dependent heat capacity parameters rather than random constant values. In Electrical Engineering, parameters are grounded in fundamental constants (e.g., $c_{0}$, $\epsilon_{0}$) and real-world materials (e.g., Polyethylene), with signal properties procedurally generated to explicitly test concepts like aliasing. In Mechanical Engineering, problems co-select consistent properties for real materials (e.g., 6061-T6 Aluminum) and are context-aware, such as using Mercury for manometer problems. A comprehensive list of these parameterization examples is provided in Appendix E, and the authoritative data sources used are listed in Appendix F.
Furthermore, across all templates, these already constrained parameters are subjected to multi-layer sanity checks. For example, a chemical engineering template verifies that a final concentration is less than the initial one. A mechanical engineering problem ensures a calculated stress does not exceed the material’s yield strength. Similarly, an electrical engineering template confirms a wave’s propagation speed does not exceed the speed of light. These checks validate that the final combination of values results in a solvable and physically meaningful scenario.
4.5 AI-Assisted Quality Assurance
To ensure the correctness and scalability of the benchmark, we developed an AI-Assisted Quality Assurance workflow. While traditional unit tests are effective, they proved impractical for our generative framework due to the high number of templates and the complexity of verifying physically plausible outputs. Our quality assurance workflow automates this process by employing an LLM as a first-pass Peer Reviewer to validate new problem templates before their inclusion in the final benchmark.
The core of this workflow is a programmatic validation script that prepares a comprehensive validation package for the LLM. For each template, this package includes not only the template’s full Python source code but also three unique instances (question-solution pairs) generated by it. This context is crucial, as it allows the LLM to assess both the static logic of the code and the dynamic quality of its output. We then pass this package to the LLM via a carefully engineered prompt. As detailed in Appendix G, the prompt instructs the LLM to assume the persona of an expert engineering professor and to evaluate the template against a multi-axis rubric provided in Appendix H.
Upon receiving the results, our system programmatically parses and evaluates the scores. A template, $T$, is automatically approved if and only if the score assigned to it for every core category ($c_{i}$) in our rubric meets or exceeds a predefined threshold, $\theta$. We formalize this approval condition as:
$\forall c_{i}\in C,\text{Score}(T,c_{i})\geq\theta$
We established a high threshold of $\theta=4$ based on empirical testing against a control set of known-good and deliberately flawed templates (Examples are provided in Appendix L). Any template failing to meet this criterion is automatically flagged for human review. This threshold demonstrated an optimal balance between maximizing automation and reliably catching potential errors.
In practice, this automated review process proved highly effective. Out of the 90 total templates, only 5 (5.55%) were flagged by the LLM and required subsequent human review. [^1] This human review was performed by the authors, who possess relevant engineering domain expertise. While we did not calculate formal inter-rater reliability between the LLM and human reviewers for this quality assurance step, we manually inspected and corrected the flagged templates.
4.6 Symbolic Generation and Combinatorial Diversity
A core feature of EngChain is its programmatic generation, designed to create a problem space so vast that it effectively prevents evaluation through memorization. Rather than relying on a static set of problems, EngChain is built upon symbolic templates where parameters are systematically randomized within domain-aware, physical constraints.
This design provides strong resistance to training data contamination, as it makes it statistically improbable that an LLM has encountered any specific problem instance. Consequently, successful performance on EngChain is stronger evidence of a model’s ability to generalize its reasoning capabilities. The model is forced to solve each problem from first principles rather than relying on pattern matching or rote recall of seen solutions.
4.7 Difficulty Scaling
Problem difficulty in EngChain is not a subjective label but a calibrated measure determined by systematically varying three primary factors:
  • Conceptual Complexity refers to the number and integration of engineering principles required. An Easy problem typically tests a single, isolated concept like Ohm’s Law. An Intermediate problem requires the combination of two or three concepts, such as applying the principle of superposition to calculate a net electric field from multiple point charges. An Advanced problem demands the integration of multiple complex ideas, such as using numerical integration on Levenspiel plot data to determine a reactor volume.
  • Mathematical Sophistication describes the nature of the required calculations. Easy problems involve direct algebraic substitution. Intermediate problems may require basic calculus or vector addition. Advanced problems escalate to solving differential equations or applying numerical methods, such as the trapezoidal rule for integration.
  • Procedural Depth measures the length and interdependence of the reasoning chain. An Easy problem might be solvable in one or two steps. In contrast, an Advanced problem involves a long sequence of calculations where the output of each step is a critical input for the next, testing a model’s ability to maintain a coherent and accurate computational trace.
This multi-axis framework for difficulty allows EngChain to move beyond a simple easy/hard dichotomy and provides a more fine-grained analysis of an LLM’s reasoning capabilities.
5 Evaluation Framework
For a verifiable assessment of the entire reasoning process, we compute several key metrics by comparing a model’s generated solution against the ground-truth reasoning trace. The evaluation is multifaceted. It jointly evaluates final answer accuracy, the fluency and semantic quality of the generated solution, and the correctness of the multi-step reasoning chain.
5.1 Final Answer Accuracy
It is determined by comparing the model’s predicted numerical answer ($A_{pred}$) to the ground-truth answer ($A_{gt}$). A response is considered correct only if its relative error is less than a strict tolerance ($\epsilon_{final}$) of 1%, a standard appropriate for engineering precision. The calculation is formalized as follows:
$\text{Relative Error}=\left|\frac{A_{gt}-A_{pred}}{A_{gt}}\right|<\epsilon_{final}$
5.2 Textual Quality
We assess the textual quality of a generated solution, evaluating its fluency and semantic fidelity against the ground-truth text using a combination of semantic and lexical similarity scores using the most common reference based evaluation metrics (Xie et al., 2023b):
  • BERTScore (Zhang et al., 2020) measures deep semantic similarity by aligning contextual embeddings of tokens from the generated solution and the ground-truth solution. It computes precision, recall, and an F1-score to provide a robust measure of semantic equivalence, capturing nuances that lexical overlap metrics may miss.
  • ROUGE-2 (Lin, 2004) quantifies the lexical overlap by calculating the F1-score based on the number of matching adjacent word pairs (bigrams) between the generated and ground-truth texts.
Together, these scores provide a complementary assessment that reflects the model’s ability to produce solutions using appropriate and correctly structured domain-specific terminology.
5.3 Multi-Step Reasoning Verification
This is the core of our evaluation framework which assesses the entire logical chain of a model’s solution. This process moves beyond final answer accuracy by comparing each predicted reasoning step ($P_{j}$) against each ground-truth step ($GT_{i}$). A step is considered valid only if it satisfies two criteria: semantic similarity and numerical correctness.
To formalize this, we compute a combined similarity matrix, $S_{comb}$, by taking the element-wise product (Hadamard product $\circ$) of a semantic similarity matrix ($S_{sem}$) and a numerical correctness matrix ($S_{num}$).
$S_{comb}=S_{sem}\circ S_{num}$
  • Semantic Similarity ($S_{sem}$) is computed using a powerful Cross-Encoder model (cross-encoder/stsb-roberta-large). This model processes each ground-truth and predicted step as a single contextual pair, yielding a high-fidelity similarity score that captures nuanced semantic relationships more effectively than bi-encoder models like the initially used SentenceTransformer.
  • Numerical Correctness ($S_{num}$) is a binary matrix where an entry is 1 if the intermediate numerical answer in a predicted step matches the corresponding ground-truth step within a 2% relative error tolerance, and 0 otherwise.
From this combined matrix, we determine if a step is “matched” by checking if its maximum similarity score exceeds a predefined threshold, $\tau$. We then calculate recall and precision to evaluate the reasoning chain. Recall measures the fraction of ground-truth steps the model successfully reproduced:
$\text{Recall}=\frac{1}{|GT|}\sum_{i=1}^{|GT|}\mathbb{I}\left(\max_{j}S_{comb}[i,j]>\tau\right)$
Precision measures the fraction of the model’s generated steps that were valid and relevant:
$\text{Precision}=\frac{1}{|P|}\sum_{j=1}^{|P|}\mathbb{I}\left(\max_{i}S_{comb}[i,j]>\tau\right)$
Finally, our primary metric for reasoning quality, the Step F1-Score, is calculated as the harmonic mean of the step-level recall and precision defined above, providing a single, robust measure of procedural correctness.
$\text{Step F1-Score}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
5.4 Automated Error Analysis with LLM-as-a-Judge
While the Step F1-Score provides a sound quantitative measure of a model’s procedural correctness, it has a significant limitation: it can identify that a reasoning chain is flawed, but it cannot explain why. For instance, as will be discussed later, our preliminary results revealed models that achieved high final-answer accuracy despite very low reasoning scores, a classic case of being “right for the wrong reasons.” This made us realize that to diagnose these failures and determine their cause, a deeper, more qualitative analysis was required. Hence, to move beyond quantitative scores and introduce a layer of explainability, we propose a novel Automated Error Analysis system that employs a large language model as an expert evaluator, or an LLM-as-a-Judge.
This system is triggered for any ground-truth step ($GT_{i}$) that fails to find a satisfactory match among the predicted steps. A step is flagged as a “failure” if its maximum combined similarity score falls below an empirically determined threshold, $\tau=0.7$. This condition is formalized as:
$\max_{j}S_{comb}[i,j]<\tau$
Figure 5: Automated error analysis with the LLM-as-a-Judge workflow.
The overall workflow of this automated analysis process is illustrated in Figure 5. At its core, the process is rooted in carefully designed prompt engineering, where the judge model is given a system prompt defining its persona as an expert engineering professor and a set of critical rules for logical consistency. The input to the judge is a structured prompt containing the full problem context, the ground-truth step, and the model’s incorrect step. Providing this complete context is crucial for the judge to accurately classify the failure into one of the five distinct categories, which are defined in Appendix I. To ensure a reliably parsable output, the judge is instructed to return only a structured JSON object containing its classification and a detailed and scalable diagnosis of the model’s specific failure modes. The full system prompt used for the automated error analysis is provided in Appendix J.
6 Experiments
6.1 Evaluation Model Suite
To provide a comprehensive evaluation of engineering reasoning capabilities, we evaluate models from a diverse suite. The results presented in Section 7 focus on 11 distinct models from the first category. While a full evaluation of the other categories is planned for future work, we group the complete suite of models into the following four-way taxonomy:
  1. 1.
    Frontier proprietary models, which represent the current upper bound of performance. This includes models such as GPT-{5, 4.1, 5-mini, 4.1-mini} (OpenAI, 2025a, b), Claude Sonnet {4.5, 4, 3.7} (Anthropic, 2025b, c, a), Gemini 2.5 {Pro, Flash} (Comanici et al., 2025), DeepSeek {V3.2, V3.1, R1} (Liu et al., 2024; Guo et al., 2025), and Grok 4 {Heavy, Fast} (xAI, 2025).
  2. 2.
    Science/Engineering-enhanced models, such as Galactica (Taylor et al., 2022), Minerva (Lewkowycz et al., 2022), ChemLLM (Zhang et al., 2024), SciLit01 (Li et al., 2025a), and SciAgent (Gao et al., 2024), are systems pre-trained or fine-tuned on vast scientific and technical corpora to improve domain-specific knowledge and complex procedural reasoning.
  3. 3.
    Math-enhanced models, such as WizardMath (Luo et al., 2023), MetaMath (Yu et al., 2023), Mathstral (Mistral, 2024), and Qwen2.5-Math (Yang et al., 2024), are systems specifically trained on mathematical and symbolic corpora to enhance their quantitative reasoning capabilities.
  4. 4.
    General-purpose open models, such as LLaMA 3.1 (Grattafiori et al., 2024) and Qwen {2.5, 3} (Qwen, 2024, 2025), provide strong, domain-agnostic foundations that are often used to create specialized variants.
This four-way taxonomy provides a structure for comparing different model architectures and training paradigms, highlighting how model scale, domain specialization, and mathematical supervision influence symbolic engineering reasoning. The detailed configurations and model sources can be found in Appendix K.
6.2 Experimental Setup
We instantiate the benchmark by sampling 15 instances per symbolic template, each with a distinct random seed, creating a total of 1,350 test cases (90 templates $\times$ 15 instances) for each model. We evaluate all models under a unified decoding configuration: a low temperature of 0.2 and a maximum token limit of 4,096. We use a low temperature (T=0.2) to ensure deterministic, high-confidence outputs and a high token limit (4,096) to allow ample space for detailed, multi-step solutions without premature cut-offs. We use a zero-shot setup with the following standardized reasoning prompt, which is used for all inference tasks:
We use a zero-shot setup with the following standardized reasoning prompt, which is used for all inference tasks:
You are an expert engineer. Solve the following
problem by providing a detailed, structured
solution. Use the exact headings and
formatting provided below.
# Given
List all known variables and their
values with units.
# Find
State the variable(s) to be calculated.
# Formulae
Write down all necessary governing equations
before substituting any values.
# Solution
Provide a step-by-step calculation. Each step
must start on a new line and be formatted exactly
as ’**Step X:**’, where X is the step number.
Show the substitution of values into
the formulae clearly.
# Final Answer
State the final numerical result with its units
in the format: **Answer:** [value] [units]
To process the model outputs, we first segment the text into reasoning steps. We use a regular expression to identify line-start markers that match either the Step <number> format (e.g., **Step 1**) or the numbered-list format (e.g., 1.). A subsequent regex cleans these prefixes. We then extract numerical values from each step and the final answer using a three-priority heuristic: first, matching the **Answer:** tag; second, falling back to the last number following an equals sign (e.g., V = 100.0); and third, taking the last standalone number. If no overall **Answer:** tag is found, the final numerical value from the last step is used for verification.
7 Results
7.1 Overall Model Performance
The primary finding from our evaluation is a stark and universal divergence between the models’ ability to produce a correct final answer and the verifiable correctness of their reasoning process. As shown in  Table 1, while the top models achieve a final answer accuracy of approximately 63.1% (gpt-4.1-mini and gpt-5-mini), their procedural reasoning capability is critically low. The Reasoning (F1 Score), our primary metric for procedural correctness, is alarmingly poor across the board, with the top-performing model, gemini-2.5-flash, only achieving 19.32%. This demonstrates that in the vast majority of cases, all models fail to follow a sound, verifiable problem-solving methodology, confirming the “right for the wrong reasons” phenomenon at scale.
This significant reasoning gap is further masked by high textual fluency. The BERTScore, for instance, remained consistently high for all models, averaging approximately 0.87. This indicates that models excel at generating semantically appropriate, engineering-domain language, even when the underlying procedural logic is flawed or nonsensical.
Model Final Answer Accuracy $\uparrow$ Reasoning (F1 Score) $\uparrow$ BERTScore $\uparrow$ ROUGE-2 $\uparrow$ ROUGE-L $\uparrow$
Frontier Proprietary Models
Claude 3.7 56.8149.55 16.9121.76 88.031.16 29.287.65 37.097.63
Claude 4 57.7849.41 14.1618.28 87.931.33 29.558.16 37.257.91
Claude 4.5 52.2249.97 14.1919.72 87.391.09 28.108.15 35.347.58
Deepseek v3.2 61.2648.73 13.5319.54 86.541.23 25.698.96 34.579.12
Deepseek R1 59.9349.02 19.0324.82 86.481.34 25.588.34 34.498.30
Gemini 2.5 Flash 59.4149.13 19.3223.56 87.551.63 28.829.17 36.409.36
Gemini 2.5 Pro 59.9349.02 17.9721.01 88.241.41 31.828.29 37.918.49
GPT 4.1 57.7849.41 12.1315.85 86.651.24 26.749.13 34.459.12
GPT 4.1 Mini 63.1148.27 14.8419.69 86.521.14 25.198.17 33.828.34
GPT 5 60.4448.92 15.4721.79 88.191.21 28.988.30 36.558.06
GPT 5 Mini 63.1148.27 14.3020.32 87.951.10 26.637.62 34.747.52
Table 1: Overall performance of frontier models on the EngChain benchmark (N=1350) under a Zero-Shot setting. Scores represent the mean, with standard deviation shown in superscript. The highest mean score in each column is in bold.
7.2 Performance by Engineering Branch
A fine-grained analysis of the results reveals a significant performance disparity across the three engineering branches, as illustrated in Figure 6. We observe that model reasoning capabilities are not uniform and are highly sensitive to the specific scientific principles and mathematical structures of a given engineering domain.
Chemical Engineering consistently proved to be the most challenging domain for all models, yielding the lowest scores in both final answer accuracy and procedural reasoning. The reasoning collapse in this domain is particularly severe, with several models falling into single-digit Reasoning (F1 Score)s, including Deepseek v3.2 (6.90%) and GPT 5 Mini (7.58%).
Conversely, Mechanical Engineering was clearly the most solvable domain, where models achieved their highest scores. This branch produced the peak Final Answer Accuracy of the entire benchmark (Gemini 2.5 Pro at 70.22%) and the highest Reasoning (F1 Score) (Deepseek R1 at 30.11%).
Electrical Engineering demonstrated intermediate performance, with scores that consistently fell between the other two branches. Its peak Reasoning (F1 Score) of 22.49% (Gemini 2.5 Flash) and peak Final Answer Accuracy of 67.11% (GPT 4.1 Mini) were modest, clearly lagging behind the top scores in Mechanical Engineering but remaining well above the challenging baseline set by Chemical Engineering.
Figure 6: Branch-level performance comparison of the top four models ranked by mean Reasoning (F1 Score) across the three engineering branches. Chemical Engineering is consistently the most challenging domain for procedural reasoning.
7.3 Performance by Engineering Domain
Further analysis reveals that model performance is highly inconsistent across the nine distinct engineering domains, as visualized in the radar plot in Figure 7. This “spiky” performance profile, where a model can excel in one domain while failing in another, suggests that current models possess patches of specialized knowledge rather than generalized, first-principles reasoning.
The data shows several domains are exceptionally challenging for all models. Domains such as Reaction Kinetics, Mechanics of Materials, Transport Phenomena, and Electromagnetics & Waves produced critically low reasoning scores. In these areas, it was common for top-tier models to score in the single digits, with Deepseek v3.2 scoring 6.90% in Reaction Kinetics and GPT 5 Mini scoring 7.58% in Mechanics of Materials.
Conversely, the domain of Vibrations & Acoustics emerged as a clear outlier, representing the easiest domain for nearly all models. The reasoning scores here were dramatically higher than in any other domain, with Deepseek R1 achieving 52.26% and Gemini 2.5 Flash achieving 40.70%. The domain of Signals & Systems was another area of relative strength, where most models achieved some of their highest reasoning scores.
This domain-level variance underscores the need for broad, multi-domain benchmarks. We hypothesize this performance gap is linked to the underlying mathematical complexity. The high-scoring domains in our results often feature problems solvable with direct algebraic substitution or linear, discrete procedures. The most challenging domains, in contrast, frequently require models to set up and solve differential equations or manage systems of complex, interdependent variables, which appear to be critical failure points for current procedural reasoning.
Figure 7: Domain-level performance comparison of the top four models ranked by mean Reasoning (F1 Score). The plot reveals a “spiky” performance profile, highlighting that model reasoning is highly specialized and not generalized across topics.
7.4 Performance by Problem Difficulty
As expected, model performance degrades as problem complexity increases, but this trend is most pronounced in Final Answer Accuracy. Most models maintain a respectable accuracy of 60-70% on Easy and Intermediate tasks. However, performance drops sharply on Advanced problems, with nearly all models falling into the 40-50% range. For instance, gemini-2.5-pro’s accuracy falls from 68.6% on Intermediate problems to just 45.2% on Advanced.
The Reasoning (F1 Score), by contrast, tells a more complex story, as illustrated in Figure 8. Instead of a uniform decline, the reasoning scores for many top models (such as claude-3-7 and gpt-5) are lowest on Intermediate tasks before recovering slightly on Advanced tasks. A few models (e.g., gemini-2.5-flash and deepseek-reasoner) even show a slight improvement in reasoning as difficulty increases. The most critical finding, however, is that reasoning scores remain universally low (sub-21%) across all three difficulty levels. This suggests the models’ procedural reasoning failures are fundamental and not simply a function of task complexity.
Figure 8: Difficulty-level performance comparison of the top four models ranked by mean Reasoning (F1 Score). The plot shows procedural reasoning performance is critically low across all levels of complexity, not just on advanced tasks.
7.5 Qualitative Error Analysis with LLM-as-Judge
While the quantitative Reasoning (F1 Score) identifies that a reasoning chain is flawed, it cannot explain why. To diagnose the root causes of these failures, we employed our LLM-as-Judge system (described in Section 5) to perform a qualitative error analysis on Gemini 2.5 Pro. We selected this model as our representative subject because it is a strong “all-rounder,” achieving both a top-tier reasoning score (17.97%) and the first rank on all language fluency metrics (BERTScore, ROUGE-2, and ROUGE-L). The analysis was done using GPT-4o (OpenAI, 2024) as the judge or expert evaluator model.
The judge was used to analyze all 4,118 reasoning steps that were flagged by our verifier as a “mismatch” from the ground-truth solution. The results, summarized in Table 2, are striking. The single most significant finding is that the vast majority of flagged steps were not errors at all. Our judge classified 73.94% (3,045 instances) as Alternative Correct or in other words, valid reasoning paths that simply differed from the ground-truth solution. This overwhelmingly indicates that models are frequently penalized by rigid, single-path evaluations for finding a valid, creative, or different solution, a phenomenon masked by low quantitative scores.
For the remaining 1,072 instances identified as genuine errors, Conceptual Error was the dominant failure mode, accounting for 59.1% of all true errors (634 instances). This demonstrates that models are far more likely to fail because they misapply a physical principle or choose the wrong formula, rather than making simple Calculation Errors (14.3%) or Input Errors (7.6%). This finding pinpoints the primary weakness of modern LLMs not in their arithmetic, but in their application of deep, domain-specific knowledge. A full breakdown of error distributions by branch, domain, and difficulty can be found in Appendix M.
Error Category Total Count Percentage (%)
Alternative Correct 3,045 73.94%
Conceptual Error 634 15.40%
Other 204 4.95%
Calculation Error 153 3.72%
Input Error 81 1.97%
Analysis Failed 1 0.02%
Total 4,118 100.0%
Table 2: Overall distribution of error types identified by Gemini 2.5 Pro as LLM-as-Judge across all 4,118 flagged reasoning mismatches.
8 Conclusion and Future Work
We introduced EngChain, a symbolic benchmark for verifiable, multi-step engineering reasoning. Built from a taxonomy of 90 problem templates across 9 domains and 3 engineering branches, EngChain provides a verifiable evaluation of multi-step engineering problem-solving, assessing the entire reasoning chain rather than only the final answer.
Our comprehensive evaluation of frontier models reveals a stark “right for the wrong reasons” phenomenon: a universal gap between modest Final Answer Accuracy and critically low Reasoning (F1 Score)s. Our qualitative analysis further reveals that this gap is not only due to error, but that a striking majority of flagged mismatches are actually Alternative Correct solutions, which single-path verifiers would wrongly penalize. For genuine failures, we find that Conceptual Errors are the dominant failure mode, not simple calculation mistakes.
Future work will proceed in several key directions. First, we will complete our planned evaluation by benchmarking the science-enhanced, math-enhanced, and general-purpose open models to provide a holistic comparison across all model categories. Second, we will expand the LLM-as-Judge analysis to other top-performing models to create diagnostic profiles of their unique failure modes. Third, to rigorously validate our qualitative methodology, we will conduct two crucial follow-up studies. A Human Evaluation Study will be performed to establish a definitive ground-truth for our judge’s classifications. This will be complemented by a Cross-Judge Consistency study to ensure our error categorizations are not an artifact of our chosen evaluator (GPT-4o (OpenAI, 2024)). This consistency study will involve a diverse set of judges, including a model from a different leading developer, such as Claude 4 Opus (Anthropic, 2025c), and a high-capability open-source model, such as Qwen 2.5 (Qwen, 2024). Finally, upon completion of these studies, all evaluation scripts, prompts, and detailed results will be publicly released to ensure full reproducibility.
Limitations
We acknowledge several limitations in this work that we plan to address in future research.
First, a significant limitation lies in our primary quantitative metric, the Reasoning (F1 Score). By design, this metric measures alignment against a single, pre-defined ground-truth path. However, our own qualitative analysis (Section 7.5) revealed that a striking majority of steps flagged as “mismatches” were actually Alternative Correct solutions. This indicates that our quantitative F1 score, when used in isolation, significantly underestimates the true reasoning capabilities of the models and inherently penalizes valid, alternative problem-solving approaches. Future work should therefore focus on developing more advanced evaluation methods, such as semantic alignment metrics or dynamic, model-based verifiers, which can programmatically accept and validate multiple correct reasoning paths.
Second, EngChain is entirely synthetic, generated from symbolic templates. While this design ensures verifiability and contamination resistance, it does not capture the linguistic diversity or contextual ambiguity of real-world engineering. More complex, ill-defined problems often require interpreting non-textual information, such as technical diagrams, schematics, or CAD models, which are not part of our current evaluation. This can be addressed in future versions by expanding the benchmark to include multi-modal tasks and by using real-world engineering documents as seeds for semi-structured problem generation.
Third, the benchmark’s scope, while covering 90 problem types, is currently limited to three core engineering branches. It does not yet include other major fields, such as civil, aerospace, or biomedical engineering. A clear next step is to expand the EngChain taxonomy in collaboration with domain experts to cover these additional fields, ensuring a more comprehensive test of engineering knowledge.
Appendix A LLM Prompt for Domain Validation
You are an expert academic and senior curriculum designer at a top-tier engineering university (like MIT or Caltech). You have decades of experience structuring undergraduate programs for ABET accreditation.
Your task is to identify and rank the most fundamental, cornerstone domains for the three primary branches of engineering. The goal is to identify the non-overlapping, core subject areas that are prerequisites for almost all other advanced topics.
For example, for Chemical Engineering, ‘‘Reaction Kinetics" would be a fundamental domain, while a niche topic like ‘‘Polymer Rheology" would be a sub-specialization, not a cornerstone.
Please provide a ranked list of the top 5--7 most fundamental domains for each of the following three branches:
1. Chemical Engineering
2. Electrical Engineering
3. Mechanical Engineering
Return your answer only in a strict JSON format as given below.
{ ‘‘chemical_engineering": [
   ‘‘Rank 1 Domain",
   ‘‘Rank 2 Domain",
   ‘‘Rank 3 Domain",
   ‘‘Rank 4 Domain",
   ‘‘Rank 5 Domain"
‘‘electrical_engineering": [
   ‘‘Rank 1 Domain",
   ‘‘Rank 2 Domain",
   ‘‘Rank 3 Domain",
   ‘‘Rank 4 Domain",
   ‘‘Rank 5 Domain"
‘‘mechanical_engineering": [
   ‘‘Rank 1 Domain",
   ‘‘Rank 2 Domain",
   ‘‘Rank 3 Domain",
   ‘‘Rank 4 Domain",
   ‘‘Rank 5 Domain"
} Do not include any preamble, conversational text, explanations, or markdown formatting around the JSON block.
The above prompt was used to query large language models to cross-reference and validate the selection of fundamental engineering domains identified in our literature review.
Appendix B LLM Prompt for Area Validation
The following prompt was used to query large language models to identify the fundamental areas or sub-topics within each core engineering domain.
You are an expert academic and senior curriculum designer at a top-tier engineering university (like MIT or Caltech). You have decades of experience structuring undergraduate programs for ABET accreditation.
Your task is to identify and list the most fundamental, non-overlapping ‘‘areas" or ‘‘sub-topics" within the provided engineering domain.
A ‘‘fundamental area" is a major pedagogical unit or chapter within a core course on this domain. The areas you list should be distinct pillars of the subject. For example, if the domain is ‘‘Thermodynamics," a fundamental area would be ‘‘The First Law" or ‘‘Properties of Pure Fluids." A concept like ‘‘Enthalpy" would be too specific. Similarly, within a ‘‘Circuit Analysis" domain, ‘‘Kirchhoff’s Voltage Law" and ‘‘Kirchhoff’s Current Law" are foundational concepts often taught together and would be too granular to list separately.
Please provide a list of the top 3-5 most fundamental areas for the following engineering domain: [DOMAIN_NAME_HERE]
Return your answer only in a strict JSON format as given below.
{ ‘‘domain": ‘‘[DOMAIN_NAME_HERE]",
 ‘‘areas": [
   ‘‘Area 1",
   ‘‘Area 2",
   ‘‘Area 3",
   ‘‘Area 4"
Do not include any preamble, conversational text, explanations, or markdown formatting around the JSON block.
Appendix C Foundational Textbooks for Problem Selection
The core engineering principles, problem typologies, and realistic constraints for the templates developed by us were sourced from the authoritative textbooks listed in Table 3. These texts, which are standard in top-tier engineering curricula, served as the primary knowledge base to ensure EngChain is pedagogically sound and grounded in the core engineering curriculum. They were also used to validate our selection of fundamental domains and areas.
Domain Textbook
Reaction Kinetics Elements of Chemical Reaction Engineering by H. Scott Fogler
Thermodynamics Introduction to Chemical Engineering Thermodynamics by J. M. Smith, H. C. Van Ness, and M. M. Abbott
Transport Phenomena Transport Phenomena by R. Byron Bird, Warren E. Stewart, and Edwin N. Lightfoot
Electromagnetics and Waves Fundamentals of Applied Electromagnetics by F. T. Ulaby and Umberto Ravaioli
Communication Systems Digital Communications by John G. Proakis and Masoud Salehi
Signals and Systems Discrete-Time Signal Processing by Alan V. Oppenheim and Ronald W. Schafer
Mechanics of Materials Mechanics of Materials by Ferdinand P. Beer, E. Russell Johnston, Jr., and John T. DeWolf
Fluid Mechanics Fundamentals of Fluid Mechanics by Bruce R. Munson, Donald F. Young, and Theodore H. Okiishi
Vibrations and Acoustics Mechanical Vibrations by Singiresu S. Rao
Table 3: Foundational textbooks used for selecting domain-specific problems.
Appendix D LLM Prompt for Pedagogical Scoring
The following prompt was used to query large language models (specifically, gemini-2.0-flash) to evaluate the pedagogical significance of each sub-topic within its parent domain.
You are an expert academic and senior curriculum designer at a top-tier engineering university with decades of experience structuring undergraduate programs for ABET accreditation.
Your task is to evaluate the pedagogical significance of a specific engineering ‘‘area" within a broader ‘‘domain" on a scale of 1 to 5, based on the rubric below.
Rubric:
Score 5 (Cornerstone): A foundational, prerequisite topic for nearly all other concepts in the domain. Score 3 (Core Concept): A standard, important topic that builds upon cornerstone principles but is not necessarily a universal prerequisite. Score 1 (Specialized Application): A focused application or an integrative topic taught after cornerstones and core concepts are mastered. Please evaluate the following area: Domain: [DOMAIN_NAME_HERE] Area: [AREA_NAME_HERE] { ‘‘area": ‘‘[AREA_NAME_HERE]",
 ‘‘pedagogical_significance_score": <integer, a score from 1 to 5>,
 ‘‘justification": <string, a one-sentence explanation for the score>.
} Do not include any preamble, conversational text, explanations, or markdown formatting around the JSON block.
Appendix E Detailed Domain-Aware Parameterization Examples
To ensure that problems are not just mathematically solvable but are also grounded in physical and engineering realism, EngChain constrains all generated values by the principles of the domain being tested. The following provides specific examples of this methodology for each engineering branch.
E.0.1 Chemical Engineering
  • Reaction Kinetics and Stoichiometry: Instead of abstract chemicals, problems select from curated lists of common gas, liquid, and biochemical reactants like Propane, Benzene, and Glucose. More importantly, many problems are built around pre-validated, balanced chemical equations to ensure that the fundamental principles of mass conservation and stoichiometry are carefully upheld from the outset.
  • Thermodynamics: Problems involving thermodynamic calculations use a multi-layered approach to realism. For common substances, a complete and consistent set of critical properties ($T_{c},P_{c},V_{c}$) is retrieved. For reaction calorimetry, problems utilize tabulated standard heats of formation ($\Delta H_{f}^{\circ}$). Crucially, for heat transfer calculations, the model moves beyond constant specific heats by incorporating temperature-dependent heat capacity parameters ($C_{p}/R=A+BT+CT^{2}+DT^{-2}$), ensuring a high degree of physical accuracy across a range of conditions.
  • Transport Phenomena and Fluid Mechanics: To ensure realism in fluid dynamics problems, templates draw from extensive tables of fluid properties, including the density and viscosity of common liquids and gases. The benchmark also includes parameters for non-Newtonian fluids (e.g., Polymer Solutions & Melts etc), allowing it to test more complex, real-world rheological behaviors beyond simple Newtonian assumptions.
E.0.2 Electrical Engineering
  • Electromagnetics and Waves: Instead of using an arbitrary number, the phase velocity of a wave is determined by selecting a specific propagation medium from a curated list (e.g., Distilled Water, Polyethylene, Fused Silica etc). This ensures that all subsequent calculations for parameters like wavelength or wave number are physically consistent with that material’s properties. The generation process is further constrained by fundamental physical constants, including the speed of light in a vacuum ($c_{0}$) and the permittivity of free space ($\epsilon_{0}$).
  • Signals and Systems: In scenarios bridging continuous and discrete-time domains, such as sampling, the sampling frequency is intentionally set well above the Nyquist rate to model ideal conversion and isolate the effects of the discrete-time system. For purely discrete-time problems, parameters are procedurally constrained to test specific concepts. For instance, signal frequencies are generated as rational multiples of $\pi$, and downsampling problems are deliberately constructed to either cause or avoid aliasing, directly testing the understanding of frequency folding. For fundamental operations like convolution, sequences are kept short and integer-valued to focus the analysis on the computational process.
  • Digital Communications: For deterministic and random signal analysis, probabilities for discrete random variables are generated as exact rational numbers that sum perfectly to one. In digital modulation scheme problems, carrier frequencies are constrained to be integer multiples of the bit rate to simplify analysis, energy per bit ($E_{b}$) and signal-to-noise ratios ($E_{b}/N_{0}$) are generated in physically typical ranges (e.g., picojoules and realistic dB values), and modulation orders ($M$) are restricted to powers of two. Furthermore, data rates are assigned practical units such as kbps, Mbps, or Gbps, grounding the problems in contemporary engineering contexts.
E.0.3 Mechanical Engineering
  • Mechanics of Materials: Problems are grounded in the behavior of real-world materials by co-selecting properties from comprehensive tables. When a material such as “6061-T6 Aluminum” or “Carbon Fiber Reinforced Polymer” is chosen, its physically consistent Young’s Modulus ($E$), Poisson’s Ratio ($\nu$), and Shear Modulus ($G$) are used for all calculations. This ensures that problems involving stress, strain, and torsional deformation are realistic. Properties are available in both SI (GPa) and US Customary (ksi) units to reflect practical engineering work.
  • Fluid Mechanics: Realism in fluid statics problems is achieved by utilizing extensive lists of fluid and solid densities. This allows for the accurate calculation of hydrostatic pressure and buoyancy forces for a wide range of scenarios, such as a steel object submerged in seawater. The parameterization is context-aware, selecting dense fluids like Mercury for manometer problems while using lighter fluids like oil or water for pipe flow scenarios.
  • Vibrations and Acoustics: For problems involving dynamic systems, parameters such as mass, spring stiffness, and damping coefficients are randomized within ranges that represent plausible physical objects, from small mechanical components to larger structures. Additionally, the amplitudes and frequencies of harmonic forcing functions are chosen in relation to the system’s natural frequency to ensure the dynamic response is well-defined and clearly demonstrates key physical phenomena like resonance or beating.
Appendix F Authoritative Data Sources for Parameterization
To ensure the physical and engineering realism of EngChain, we built our domain-aware parameterization (discussed in Section 4.4) on authoritative data. We extracted and manually verified extensive lists of physical constants and material properties from the standard handbooks and data sources listed in Table 4.
Engineering Branch Authoritative Data Sources
Chemical Engineering Perry’s Chemical Engineers’ Handbook
NIST Chemistry WebBook
CRC Handbook of Chemistry and Physics
Electrical Engineering IEEE 100: The Authoritative Dictionary of IEEE Standards Terms
Standard Handbook for Electrical Engineers
ART-DEIT Database
Mechanical Engineering ASM Handbook, Volume 1
Marks’ Standard Handbook for Mechanical Engineers
Shigley’s Mechanical Engineering Design
Table 4: Primary data sources referenced for engineering parameterization across domains.
Appendix G LLM Prompt for AI-Assisted Quality Assurance (QA)
System Prompt
You are an expert engineering professor and a senior Python developer acting as a peer reviewer for the EngChain benchmark.
Your task is to meticulously evaluate a new problem template based on its source code and several example outputs.
Analyze the provided information and then respond ONLY with a single, valid JSON object that strictly adheres to the schema described in the user prompt. Do not add any explanatory text or markdown formatting around the JSON object.
User Prompt
Please evaluate the following engineering problem template.
1. Template Source Code:
python {template_code}
2. Generated Instances from the Template:
Instance 1:
- Question: ‘‘{q1}"
- Solution: ‘‘{s1}"
Instance 2:
- Question: ‘‘{q2}"
- Solution: ‘‘{s2}"
Instance 3:
- Question: ‘‘{q3}"
- Solution: ‘‘{s3}"
3. Evaluation Rubric & JSON Schema:
Evaluate the template based on the rubric below. The human_review_flag should be true if any score is less than 4. The explanation should be a concise, one-sentence justification for the scores and the flag.
{ ‘‘physical_plausibility_score": <integer, a score from 1--5 based on whether the problem respects the laws of physics and engineering>,
‘‘mathematical_correctness_score": <integer, a score from 1--5 based on whether the equations and calculations are accurate>,
‘‘pedagogical_clarity_score": <integer, a score from 1--5 based on whether the problem statement is clear, unambiguous, and solvable>,
‘‘confidence_score": <integer, a score from 1--5 indicating your confidence in this evaluation>,
‘‘human_review_flag": <boolean, true if the template requires human inspection, otherwise false>,
‘‘explanation": "<string, a concise, one-sentence justification for the scores and flag>"
Appendix H Multi-Axis Rubric for LLM-Based Template Validation
As part of our AI-Assisted Quality Assurance (QA) workflow, the LLM Peer Reviewer evaluates each template against a set of criteria. The multi-axis rubric used for this validation, detailing each scoring category and its description, is defined in Table 5.
Score Category Description
Physical Plausibility Does the problem scenario respect the laws of physics and engineering?
Mathematical Correctness Are the equations used correct, and is the final calculation accurate based on the inputs?
Pedagogical Clarity Is the problem statement clear, unambiguous, and solvable with the information provided?
Confidence Score How confident is the LLM in its own validation of this template?
Human Review Flag Based on the scores above, does this template require human inspection?
Explanation A concise, one-sentence justification for the scores and the flag.
Table 5: Multi-axis rubric used for LLM-based template validation. Numerical criteria are rated on a 1–5 scale, and the Human Review Flag is a binary (true/false) indicator.
Appendix I LLM-as-Judge Error Categories
During the second stage of our evaluation framework (Section 5), the LLM-as-a-Judge system performs a qualitative error diagnosis on failed reasoning steps. This provides deeper insight than a quantitative score alone by distinguishing between fundamental reasoning failures and valid alternative solutions. The five distinct error categories used for this classification are defined in Table 6.
Error Category Description
Conceptual Error The model applied the wrong scientific principle or formula (e.g., used addition instead of subtraction).
Calculation Error The model used the correct formula but made a mathematical mistake (e.g., $2\times 3=5$).
Input Error The model used the correct formula but pulled the wrong number from the problem context or a previous step.
Alternative Correct The model’s step is valid and logically sound but follows a different method or phrasing.
Other The model’s step is nonsensical, irrelevant, a hallucination, or contains only formatting errors.
Table 6: Error taxonomy used to categorize model mistakes during LLM-as-Judge evaluation.
Appendix J LLM Prompt for Automated Error Analysis (LLM-as-a-Judge)
You are an expert engineering professor acting as an automated evaluator. Your task is to analyze the ‘‘MODEL’S STEP" against the ‘‘GROUND-TRUTH STEP" and return a structured JSON object based on the rules and inputs below. CRITICAL RULES: 1. Your ‘‘explanation" MUST logically justify your chosen ‘‘error_category". Do not contradict yourself. For example, do not choose ‘‘Calculation Error" and then state that the calculation is correct. 2. If the model’s step is factually correct but takes a different path than the ground-truth, you MUST use the ‘‘Alternative Correct" category. Do not classify a correct step as ‘‘Other". 3. The final output must be only a raw JSON object. Do not include any introductory text, concluding remarks, or markdown formatting. Error Categories: ‘‘Conceptual Error": The model applied the wrong scientific principle or formula (e.g., used addition instead of subtraction). ‘‘Calculation Error": The model used the correct formula but made a mathematical mistake (e.g., 2 * 3 = 5). ‘‘Input Error": The model used the correct formula but pulled the wrong number from the problem context or a previous step. ‘‘Alternative Correct": The model’s step is valid and logically sound, but follows a different method or phrasing than the ground-truth step. ‘‘Other": The model’s step is nonsensical, irrelevant, a hallucination, or contains only formatting errors. Input for Analysis:
 [CONTEXT]: {problem_context}
 [GROUND-TRUTH STEP]: {gt_step}
 [MODEL’S STEP]: {pred_step}
OUTPUT FORMAT: You must now provide your analysis. Your entire response will be a single, raw JSON object. Adhere strictly to the following format with exactly two keys:
{"error_category": "...", "explanation": "..."}
Appendix K Evaluated Model Details
This appendix provides a comprehensive overview of all Large Language Models discussed in our experimental framework (Section 6). As outlined in the main text, our evaluation suite is structured around a four-way taxonomy to enable a structured comparison across different training paradigms. These four categories are: (1) Frontier Proprietary LLMs, (2) Science / Engineering Enhanced LLMs, (3) Math Enhanced LLMs, and (4) General Purpose Open LLMs.
Table 7 provides the complete list of models within each category, detailing their developing organization, parameter size (where public), backbone model (for fine-tuned variants), and source identifier (either an API endpoint or a HuggingFace repository path).
Model Organization Size Backbone Source
Frontier Proprietary LLMs
GPT-5 OpenAI N/A gpt-5-2025-08-07
GPT-4.1 OpenAI N/A gpt-4.1-2025-04-14
GPT-5 mini OpenAI N/A gpt-5-mini-2025-08-07
GPT-4.1 mini OpenAI N/A gpt-4.1-mini-2025-04-14
Claude Sonnet 4.5 Anthropic N/A claude-sonnet-4-5-20250929
Claude Sonnet 4 Anthropic N/A claude-sonnet-4-20250514
Claude Sonnet 3.7 Anthropic N/A claude-3-7-sonnet-20250219
Gemini 2.5 Pro Google DeepMind N/A gemini-2.5-pro-2025
Gemini 2.5 Flash Google DeepMind N/A gemini-2.5-flash-2025
DeepSeek V3.2 DeepSeek N/A deepseek-v3.2-20250929
DeepSeek V3.1 DeepSeek N/A deepseek-v3.1-20250922
DeepSeek R1 DeepSeek N/A deepseek-r1-20250120
Grok 4 Heavy xAI N/A grok-4-heavy-0709
Grok 4 Fast xAI N/A grok-4-fast-reasoning
Science / Engineering Enhanced LLMs
Galactica Meta AI 120B facebook/galactica-120b
Minerva Google DeepMind 540B deepmind/minerva
ChemLLM Shanghai AI Lab 13B meta-llama/Llama-2-13B ShanghaiAILab/ChemLLM-13B
SciLit01 TII 7B Qwen/Qwen2-7B-Instruct tiiuae/SciLit-01-7B
SciAgent OpenBioLLM Team 8B mistralai/Mistral-7B-v0.3 OpenBioLLM/SciAgent-8B
Math Enhanced LLMs
WizardMath WizardLM Team 7B mistralai/Mistral-7B-v0.1 WizardLMTeam/WizardMath-7B-V1.1
MetaMath MetaMath Project 7B EleutherAI/llemma-7b meta-math/MetaMath-7B-V1.0
Mathstral Mistral AI 7B mistralai/Mistral-7B-v0.1 mistralai/Mathstral-7B-v0.1
Qwen2.5-Math Qwen Team 7B Qwen/Qwen2.5-7B Qwen/Qwen2.5-Math-7B-Instruct
General Purpose Open LLMs
LLaMA 3.1 Meta 8B meta-llama/Llama-3.1-8B
Qwen 2.5 Alibaba Qwen Team 7B Qwen/Qwen2.5-7B-Instruct
Qwen 3 Alibaba Qwen Team 8B Qwen/Qwen3-8B
Table 7: Overview of all models evaluated, grouped by functionality: (1) frontier proprietary LLMs representing current performance upper bounds, (2) science/engineering-enhanced models trained on technical corpora, (3) math-enhanced models for quantitative reasoning, and (4) general-purpose open LLMs providing broad foundations. For proprietary models, we list the internal or API model identifiers; for open models, we include the HuggingFace source.
Appendix L AI-Assisted QA Workflow: Template Examples
This appendix provides visual evidence of the LLM Peer Reviewer’s output for an automatically approved template (provided in Figure 9) and a flagged template (provided in Figure 10), illustrating the function of our automated Quality Assurance (QA) process (Section 4.5).
Figure 9: LLM Peer Review result for an Approved Template (template_rackett_equation_volume). The template, designed for Chemical Engineering Thermodynamics, receives the maximum score of 5/5 across all core categories (Physical Plausibility, Mathematical Correctness, and Pedagogical Clarity), indicating that the problem structure is robust and ready for benchmark inclusion.
Figure 10: LLM Peer Review result for a Flagged Template (template_euclidean_distance_binary). The model detected a fundamental flaw, assigning a score of 3/5 for both Plausibility and Correctness. The explicit explanation flags a “calculation error that sets sqrt(Eb) to zero,” which immediately forces the template out of the automated pipeline and into human review for manual correction.
Appendix M Detailed LLM-as-Judge Error Analysis
This appendix provides the full, fine-grained breakdown of the qualitative error analysis performed on Gemini 2.5 Pro, using GPT-4o as the judge. The following tables detail the distribution of all 4,118 reasoning steps that were flagged by the quantitative verifier as a mismatch from the ground-truth solution. Table 8 breaks down the errors by engineering branch,  Table 9 shows the distribution by problem difficulty, and Figure 11 provides a proportional breakdown of error types across all nine engineering domains.
Branch Alternative Correct Conceptual Error Calculation Error Input Error Other Total
Chemical 1353 206 67 17 43 1686
Electrical 827 221 23 24 94 1189
Mechanical 865 207 63 40 67 1242
Total 3045 634 153 81 204 4117
Table 8: Branch-level Error Analysis for Gemini 2.5 Pro. This table shows the distribution of 4,117 flagged reasoning mismatches, categorized by the judge (GPT-4o), across the three main engineering branches.
Difficulty Alternative Correct Conceptual Error Calculation Error Input Error Other Total
Easy 1088 162 52 15 61 1378
Intermediate 1178 204 65 51 73 1571
Advanced 779 268 36 15 70 1168
Total 3045 634 153 81 204 4117
Table 9: Difficulty-level Error Analysis for Gemini 2.5 Pro. This table shows the distribution of 4,117 flagged reasoning mismatches, categorized by the judge (GPT-4o), across the three problem difficulty levels.
Figure 11: Domain-level Error Analysis for Gemini 2.5 Pro. This 100% stacked bar chart illustrates the proportional distribution of error types across all nine engineering domains. This visualization highlights how the ratio of error types changes by domain; for instance, Fluid Mechanics and Signals & Systems show a much larger proportion of Conceptual Errors and Other errors compared to Thermodynamics or Reaction Kinetics, where Alternative Correct solutions dominate.

### v2

@@
-EngChain: A Symbolic Benchmark for Verifiable Multi-Step Reasoning in Engineering
+EngTrace: A Symbolic Benchmark for Verifiable Process Supervision of Engineering Reasoning
 Abstract
-Large Language Models (LLMs) are increasingly being applied to specialized, high-stakes domains like engineering, which demands rigorous evaluation of their complex reasoning capabilities.
-While current benchmarks assess language understanding, factual recall, mathematics or code generation, none capture the integrative reasoning central to engineering where scientific principles, quantitative modeling and practical constraints must converge.
-To address this gap, we introduce EngChain, a benchmark for verifiable multi-step engineering problem-solving.
-EngChain contains 90 problems spanning three engineering branches, organized into 9 domains and 20 distinct areas.
-The problems are generated from symbolic templates with a high degree of randomization to ensure diversity and eliminate the risk of contamination.
-With this benchmark, we move beyond final answer accuracy with a two-stage evaluation: we first quantitatively verify the numerical and semantic validity of each reasoning step and then introduce LLM-as-a-Judge, an automated system to qualitatively categorize the identified reasoning errors.
-EngChain: A Symbolic Benchmark for Verifiable Multi-Step Reasoning in Engineering
-Ayesha Gull1 Muhammad Usman Safder1 Rania Elbadry2 Preslav Nakov2 Zhuohan Xie2 1Namal University 2MBZUAI ProjectCodeApp
-**footnotetext: Equal contribution.
+Large Language Models (LLMs) are increasingly entering specialized, safety-critical engineering workflows governed by strict quantitative standards and immutable physical laws, making rigorous evaluation of their reasoning capabilities imperative.
+However, existing benchmarks such as MMLU, MATH, and HumanEval assess isolated cognitive skills, failing to capture the physically grounded reasoning central to engineering, where scientific principles, quantitative modeling, and practical constraints must converge.
+To enable verifiable process supervision in engineering, we introduce EngTrace, a symbolic benchmark comprising 90 templates across three major engineering branches, nine core domains and 20 distinct areas.
+Through domain-aware parameterization, we generate 1,350 unique, contamination-resistant test cases to stress-test generalization.
+Moving beyond outcome matching, we introduce a verifiable two-stage evaluation framework that uses a tiered protocol to validate intermediate reasoning traces alongside final answers through automated procedural checks and a heterogeneous AI Tribunal.
+Our evaluation of 24 leading LLMs reveals a distinct trade-off between numeric precision and trace fidelity, identifying a complexity cliff where abstract mathematical pre-training fails to translate into the integrative reasoning required for advanced engineering tasks.
+EngTrace: A Symbolic Benchmark
+for Verifiable Process Supervision of Engineering Reasoning
+Ayesha Gull1 Muhammad Usman Safder1 Rania Hossam2 Fan Zhang3 Veselin Stoyanov2 Preslav Nakov2 Zhuohan Xie2 1Namal University 2MBZUAI 3The University of Tokyo Project Code App
 1 Introduction
-As large language models (LLMs) transition from general-purpose language tasks to specialized, high-stakes domains such as scientific discovery, industrial design, and engineering analysis, the need for rigorous and meaningful evaluation has become significant (Zhao et al., 2023; Xie et al., 2023a).
-The progress of the field is gated by the quality of its benchmarks: without tests that capture practical, physically grounded reasoning, it is difficult to separate genuine capability from sophisticated mimicry.
-To meet this challenge, we introduce EngChain, a new benchmark for verifiable, multi-step engineering problem-solving.
-Our benchmark is built on a foundation of symbolic templates, each capable of programmatically generating a vast number of unique, verifiable problems.
-Figure 1 provides a concrete example from our benchmark, illustrating how a single template generates a complete problem instance along with its full chain-of-thought solution.
-Figure 1: An example from EngChain (CSTR Volume Calculation). A symbolic template (1) generates a unique problem instance (2) and its verifiable, step-by-step chain-of-thought solution (3).
-This design directly targets two key problems in current LLM evaluation.
-The first is “benchmark saturation” (Ott et al., 2022), where models rapidly achieve superhuman performance on static benchmarks like GLUE (Wang et al., 2019) and even their more complex successors such as BBH (Kazemi et al., 2025).
-The second is a structural limitation: most benchmarks evaluate critical skills in disciplinary silos.
+With the expansion of LLMs into high-stakes domains such as engineering, the need for their comprehensive evaluation has become paramount (Zhao et al., 2023; Xie et al., 2023a).
+Progress is currently gated by benchmark quality; without tests that enforce physically grounded reasoning, distinguishing genuine capability from sophisticated mimicry remains a challenge.
+Figure 1: An example from EngTrace demonstrating CSTR Volume Calculation.
+A symbolic template generates a unique problem instance paired with a gold-standard reasoning trace to enable verifiable process supervision.
+Extending the symbolic-template paradigm pioneered in mathematics (Mirzadeh et al., 2024) and recently in finance (Xie et al., 2025) to the interdependent physical systems and specialized design constraints of engineering, we introduce EngTrace, a new benchmark designed to enable verifiable process supervision of reasoning in engineering problem-solving.
+It is built on a foundation of symbolic templates, each capable of programmatically generating a vast number of unique problem instances.
+Figure 1 illustrates a Continuous Stirred-Tank Reactor (CSTR) volume calculation example from our benchmark.
+Unlike static benchmarks, this generated instance requires synthesizing interdependent stoichiometry and reaction kinetics variables to derive the reactor size required for a specific outlet flow rate, yielding a complex reasoning trace fully verifiable against the template’s ground truth.
+EngTrace addresses two critical gaps that compromise the validity of current LLM evaluation paradigms for engineering: benchmark saturation and disciplinary fragmentation.
+First, static benchmarks like Glue (Wang et al., 2019) and BBH (Kazemi et al., 2025) are increasingly prone to model saturation (Ott et al., 2022) and data contamination (Deng et al., 2024).
+EngTrace addresses these risks through symbolic templates and domain-aware parameterization, generating unique, physically grounded problems that prevent rote memorization.
+Second, most benchmarks evaluate critical skills in disciplinary silos.
 General knowledge benchmarks (e.g., MMLU (Hendrycks et al., 2020)) test broad factual recall, while specialized benchmarks (e.g., Math (Hendrycks et al., 2021b) or HumanEval (Chen et al., 2021)) assess abstract logic and algorithmic translation, respectively.
 This fragmentation is limiting for engineering which is fundamentally an integrative discipline that requires the synthesis of these isolated skills, combining scientific principles, mathematical modeling, and practical constraints (e.g., Moaveni, 2019; Dym et al., 2012).
-This holistic, procedural reasoning is not captured by existing benchmarks (e.g., Felten et al., 2025; Li et al., 2025b), which predominantly focus on final-answer accuracy.
-In engineering, where a flawed process can lead to catastrophic failure, this evaluation gap is critical.
-To address this gap, EngChain uses symbolic templates to produce integrative reasoning problems and a two-stage evaluation via LLM-as-a-Judge that verifies the full reasoning chain rather than just the final answer.
-Verifiable multi-step reasoning is fundamental to building trustworthy AI systems across all domains.
-Concurrent with our work, FinChain (Xie et al., 2025) introduced a symbolic benchmark for verifiable chain-of-thought financial reasoning.
-However, the two benchmarks address fundamentally different reasoning types.
-FinChain evaluates reasoning based on economic principles, whereas EngChain evaluates reasoning grounded in the physical laws and practical constraints of engineering.
-Moreover, FinChain uses a global alignment metric to score the entire reasoning path, while EngChain employs a two-stage evaluation that provides both a quantitative Step F1-Score for procedural correctness and a qualitative, diagnostic error analysis using our LLM-as-a-Judge system.
-In summary, our contributions are threefold:
-  1. 1.
-    We introduce EngChain, a new symbolic benchmark for verifiable, multi-step engineering problem-solving with 90 problem templates across nine domains.
-  2. 2.
-    We propose a rigorous two-stage evaluation framework, including our qualitative LLM-as-a-Judge system, that moves beyond final answer accuracy to diagnose reasoning failures.
-  3. 3.
-    We present a comprehensive analysis of 11 frontier LLMs, revealing a critical “right for the wrong reasons" gap and identifying that the majority of reasoning failures are due to Conceptual Errors, not calculation slips.
+EngTrace addresses this structural limitation through integrative reasoning templates.
+Unlike existing benchmarks (e.g., Zhou et al., 2025; Mudur et al., 2025) that rely solely on outcome matching, our templates demand holistic, procedural reasoning.
+This is a crucial requirement in engineering, where a flawed process can lead to catastrophic failure (Leveson, 2012).
+In summary, our contributions are threefold: (1) We introduce EngTrace, a gold-standard symbolic benchmark for verifiable process supervision, comprising 90 problem templates spanning three major engineering branches, nine core domains, and 20 distinct areas.
+(2) We propose a verifiable two-stage evaluation framework that employs a tiered verification protocol utilizing automated procedural checks and a multi-model AI tribunal to assess reasoning process quality in addition to final answer accuracy.
+(3) We evaluate 24 models across various categories, revealing critical weaknesses in reasoning quality, with non-frontier models failing to generalize beyond basic patterns and performing poorly under complex engineering constraints.
 2 Related Work
-2.1 Benchmarks in Mathematics and Coding
-A large body of work on multi-step reasoning benchmarks has focused on foundational, yet isolated, cognitive skills, particularly in mathematics and coding.
-In mathematical reasoning, benchmarks ranging from grade-school problems (GSM8K)  (Cobbe et al., 2021) to high-school and college-level competition problems (MATH)  (Hendrycks et al., 2021b) establish a clear standard for evaluating abstract logical deduction.
-This is further refined by HARDMATH  (Fan et al., 2024) for applied mathematics and GSM-Symbolic  (Mirzadeh et al., 2024) for programming and symbolic tasks.
-Similarly, benchmarks for coding and software engineering evaluate models across different levels of complexity.
-HumanEval  (Chen et al., 2021) focuses on translating natural language instructions into small, correct code snippets.
-MBPP  (Austin et al., 2021) and its derivative MBUPP extend this by testing models on a broader set of programming problems with varying difficulty.
-SWE-Bench  (Jimenez et al., 2024) addresses complex, real-world codebases, measuring a model’s ability to solve practical software engineering challenges.
-While these benchmarks are essential for capturing foundational skills, their focus on isolated abilities limits their usefulness for engineering.
-A model can excel at solving abstract equations or writing pure algorithms yet still be unable to apply those skills to a problem constrained by physical laws and material properties.
-2.2 Benchmarks in the Physical Sciences
-Recognizing this limitation, later benchmarks focus on applied reasoning within the physical sciences.
-UGPhysics  (Xu et al., 2025) provides broad coverage and tests how models apply mathematical reasoning to physical laws.
-Building on this, PhysReason  (Zhang et al., 2025a) emphasizes process-oriented problem solving, guiding models through step-by-step solutions.
-ABench-Physics  (Zhang et al., 2025b) introduces highly challenging tasks to evaluate advanced capabilities, while PHYBench  (Qiu et al., 2025) targets symbolic computation.
-NEWTON  (Wang et al., 2023) assesses qualitative understanding, complementing the more quantitative approaches.
-Other benchmarks, such as LLM-SRBench  (Shojaee et al., 2025), focus specifically on quantitative scientific modeling.
-Together, these benchmarks show how models apply mathematical reasoning to well-defined physical laws, marking an important advance towards assessing real-world problem-solving.
-However, they remain largely confined to physics and have not been widely adopted for evaluating models in real-world engineering practices.
-2.3 Benchmarks in Engineering
-Current approaches to evaluating engineering reasoning are fractured, split between benchmarks that are overly broad and those that are deep but methodologically limited.
-On one hand, generalist benchmarks like MMLU  (Hendrycks et al., 2021a) and BIG-Bench  (Luo et al., 2024) include engineering topics, but their reliance on multiple-choice, knowledge-based questions primarily tests factual recall rather than the deep, procedural reasoning required for practical engineering problem-solving.
-On the other hand, more specialized engineering benchmarks, while deeper in scope, either remain confined to a single scientific discipline or lack a scalable, deterministic method for verifying the reasoning process itself.
-For example, EngiBench  (Felten et al., 2025) provides broad coverage but uses rubric-based scoring for its more complex problems, whereas FEA-Bench  (Li et al., 2025b) evaluates a model’s ability to use professional simulation software, yet this measures tool proficiency rather than fundamental reasoning.
-The preceding review makes it clear that the field lacks a benchmark that captures the integrative, cross-disciplinary, and verifiable reasoning central to engineering.
-To address this critical gap, we introduce EngChain, a benchmark for verifiable, multi-step engineering reasoning.
-It uses a two-stage evaluation that combines quantitative verification of each intermediate step with a qualitative error diagnosis, covering 90 distinct problem types across nine domains and three engineering branches.
-3 EngChain Taxonomy
-EngChain covers a wide range of core engineering problems, focusing on domains that are central to real-world practice, which we show in Figure 2.
-In Chemical Engineering, it targets the three pillars of chemical process design and analysis.
-Kinetics problems test the modeling of reaction rates and mechanisms, requiring reasoning about how systems evolve over time.
-Thermodynamics problems center on energy balances and phase behavior, capturing the constraints that govern feasibility and efficiency.
-Transport phenomena problems address momentum, heat, and mass transfer, linking microscopic principles to large-scale process behavior.
-In Electrical Engineering, it covers the full spectrum from fundamental physics to applied systems.
-At the foundation, it includes electromagnetics, which governs the behavior of fields and waves.
-Building on this, it addresses signals and systems, focusing on representation, transformation, and filtering of information.
-Finally, it extends to communications, where these principles are applied to the reliable transmission and processing of signals in real-world contexts.
-In Mechanical Engineering, it spans three core domains.
-Fluid mechanics problems address laminar and turbulent flow, testing how models capture momentum and energy transfer in fluids.
-Mechanics of materials emphasizes stress–strain relationships, deformation, and failure analysis, reflecting how structures respond to applied loads.
-Vibrations problems focus on natural frequencies and damping, capturing the dynamic stability of mechanical systems.
-Together, these areas test how forces, motion, and stability interact in physical systems.
-Figure 2: The taxonomy of EngChain.
-By integrating these principle subfields across the main branches of engineering, EngChain provides a comprehensive foundation of the analytical and procedural reasoning skills essential for engineering problem-solving.
-4 EngChain Design and Methodology
-To ensure the rigor, scalability, and long-term viability of the benchmark, we adopted a programmatic, template-based methodology.
-Unlike static datasets, which are finite and susceptible to being inadvertently included in the training corpora of future models, a programmatic approach offers several key advantages.
-It provides virtually limitless scalability and diversity, as a single template can generate thousands of unique problem instances.
-This design also ensures strong resistance to training data contamination, as it tests for generalizable reasoning on novel problems rather than the memorization of previously seen solutions.
-4.1 Content Selection Framework
-The selection of content for EngChain followed a principled, multi-stage methodology designed to ensure the benchmark is both representative of a formal engineering curriculum and grounded in foundational, real-world principles.
-The process involved a hierarchical selection of domains, areas, and finally, the specific problems to be templatized.
-4.1.1 Domain and Area Selection
-The nine core domains that form the backbone of EngChain were identified using a two-pronged approach.
-First, a thorough literature review was conducted, surveying established engineering taxonomies and curricular standards set by leading professional and accreditation bodies.
-This involved analyzing the program criteria from organizations like the Accreditation Board for Engineering and Technology (ABET, 2024), which mandates core competencies in areas like thermodynamics and circuit analysis for accredited degrees.
-We also reviewed the educational frameworks of major professional societies, such as the American Society of Mechanical Engineers (ASME, 2025), the Institute of Electrical and Electronics Engineers (IEEE, 2025), and the American Institute of Chemical Engineers (AIChE, 2025), to ensure our selected domains represent a consensus on the foundational knowledge required for professional practice.
-Second, these findings were cross-referenced by querying an LLM (gemini-2.0-flash).
-The model was prompted to identify and rank the most fundamental, cornerstone domains within each of the three primary engineering branches.
-The full prompt used for this domain-level validation is provided in Appendix A.
-The final set of nine domains represents a strong consensus between traditional academic structures and the knowledge synthesized by modern AI.
-A similar two-pronged methodology was then used to identify the most critical areas within each domain: first, by extracting the core pedagogical units from the authoritative textbooks, and second, by cross-referencing this list using a LLM (gemini-2.0-flash).
-The full prompt used for this area-level validation is provided in Appendix B.
-4.1.2 Template Selection
-Once the domains and areas were defined, we developed the 90 problem templates.
-The core engineering principles, problem typologies, and typical constraints for these templates were sourced from the authoritative and widely-adopted textbooks in each field.
-This textbook-driven approach ensures that every problem in EngChain is pedagogically sound and central to engineering practice.
-A complete list of these foundational texts is provided in Appendix C.
-Crucially, the number of templates developed for each area was not uniform.
-Instead, it was guided by a Pedagogical Significance Score assigned to each area.
-This score was programmatically determined by querying an expert-persona LLM (gemini-2.0-flash) to rate each area’s centrality to the curriculum on a 1-5 scale, as detailed in Appendix D.
-Foundational concepts with higher scores received more extensive coverage.
-For example, within the Reaction Kinetics domain, the area of Mole Balances (Significance Score: 5) was covered by five templates, while the more specialized application of Levenspiel Plot Interpretation (Significance Score: 1) was covered by a single template.
-Therefore, the varying number of templates per area is a direct and intentional outcome of our methodology.
-This ensures that EngChain’s structure reflects the hierarchical and foundational nature of engineering knowledge as it is taught and practiced.
+Benchmarks in Mathematics and Coding.
+Extensive work has established baselines for isolated cognitive skills.
+In mathematics, benchmarks ranging from GSM8K (Cobbe et al., 2021) and Math (Hendrycks et al., 2021b) to HardMath (Fan et al., 2024), Putnam-Axiom (Gulati et al., 2024), and GSM-Symbolic (Mirzadeh et al., 2024) rigorously test abstract logical deduction and reasoning stability.
+Similarly, coding evaluations span from function-level generation in HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) to real-world software engineering in SWE-Bench (Jimenez et al., 2024).
+While foundational, these benchmarks lack engineering utility, as proficiency in abstract logic does not translate to reliable reasoning under physical constraints (Heesch et al., 2025).
+Recently, FinChain (Xie et al., 2025) utilized symbolic templates to evaluate financial reasoning.
+While this demonstrated the utility of verifiable workflows, EngTrace extends the approach to physically grounded simulation, where reasoning traces are governed by immutable conservation laws and interdependent system dynamics.
+Benchmarks in the Physical Sciences.
+To bridge the gap between abstract mathematics and physical reality, recent benchmarks have pivoted to the physical sciences.
+While UGPhysics (Xu et al., 2025) and PhysReason (Zhang et al., 2025a) effectively test the application of reasoning to natural laws, others target advanced symbolic (Zhang et al., 2025b; Qiu et al., 2025) or qualitative (Wang et al., 2023; Shojaee et al., 2025) understanding.
+Despite this progress, these benchmarks remain confined to theoretical physics and lack the integrative context of engineering.
+Benchmarks in Engineering.
+This fragmentation extends to engineering evaluation.
+While generalist suites like MMLU (Hendrycks et al., 2021a),BIG-Bench (Luo et al., 2024), and SuperGPQA (Du et al., 2025) include engineering topics, their reliance on knowledge-based multiple-choice questions restricts assessment to factual recall rather than procedural fidelity.
+Conversely, specialized benchmarks offer depth but are constrained by narrow disciplinary scopes or a lack of scalable reasoning verification.
+For instance, specialized benchmarks like TransportBench, APBench, and EEE-Bench are siloed within narrow sub-disciplines (Syed et al., 2024; Chen et al., 2025; Cheng et al., 2024; Li et al., 2025; Liu and others, 2025).
+Others focus on tangential skills, such as design generation (Guo et al., 2025b) or software proficiency (Mudur et al., 2025), rather than fundamental reasoning, while broader efforts like EngiBench (Zhou et al., 2025) rely on subjective rubric-based scoring.
+Evidently, existing benchmarks remain limited to outcome matching and fail to validate the physically grounded reasoning required for engineering.
+EngTrace addresses this limitation by introducing a symbolic benchmark that pairs unique problem instances with gold-standard reasoning traces to enable verifiable process supervision.
+3 EngTrace Design and Methodology
+3.1 Taxonomy and Content Selection
+The taxonomy of EngTrace, illustrated in Figure 2, systematically covers nine core domains and 20 distinct areas across three engineering branches.
+This structure is not arbitrarily defined but follows a principled, two-pronged methodology to ensure it mirrors formal engineering curricula.
+Figure 2: EngTrace’s Hierarchical Taxonomy spanning nine domains across three branches: process equilibrium and dynamics (Chemical), field theory and signal transmission (Electrical), and physical mechanics and material properties (Mechanical).
+Domain and Area Selection.
+To identify core domains and critical areas, we synthesized engineering curricular standards from ABET (ABET, 2024) and major professional societies (ASME (ASME, 2025), IEEE (IEEE, 2025), AIChE (AIChE, 2025)).
+We validated these foundational units by cross-referencing them with an expert-persona LLM to ensure alignment between traditional pedagogy and modern AI capabilities (prompts provided in Appendix A and Appendix B).
+Template Selection.
+We developed 90 symbolic templates by sourcing core engineering principles, problem typologies, and typical constraints from foundational textbooks (e.g., Fogler, 2020; Ulaby and Ravaioli, 2015), ensuring pedagogical grounding (see Appendix C for the complete list of source texts).
+The template distribution per area is non-uniform and guided by a Pedagogical Significance Score (1–5) derived through an expert-persona LLM (detailed prompt provided in Appendix D).
+This weighting ensures hierarchical fidelity: for instance, within the Reaction Kinetics domain, the core area of Mole Balances (Score: 5) includes five templates, while the specialized Levenspiel Plot Interpretation (Score: 1) is represented by a single template.
 3.2 Template Generation Pipeline
-The template generation process follows a systematic pipeline as shown in Figure 3.
-First, we generate domain-aware parameters are generated.
-Next, the core engineering computations are performed.
-These results are then thoroughly validated to ensure they are physically plausible.
-Finally, the validated content is assembled into a natural language question and its corresponding step-by-step solution.
-Figure 3: Our template generation pipeline.
-We format the output strings to be both human-readable, like a textbook example, and machine-parsable for automated evaluation.
-After template functions are defined, we generate the problem instances by repeatedly calling each template, with every run using a unique random seed for reproducibility.
-4.3 Template Structure and Outputs
-The core of our benchmark is the template function, a standardized Python function that serves as the blueprint for a specific type of engineering problem.
-To ensure consistency, every template sticks to a simple, standardized signature i.e., it takes no arguments and returns a (question, solution) tuple.
-The question is a formatted string containing the problem statement to be presented to the LLM, whereas the solution is a formatted string containing the complete reasoning trace rather than only a numerical answer.
-A generic multi-disciplinary template structure is provided in Figure 4.
-Figure 4: Generic multi-disciplinary template structure.
-Each solution is structured like a textbook example such that the principles are stated, values are substituted, and intermediate steps are shown with consistent units.
-4.4 Domain-Aware Parameterization and Sanity Checks
-EngChain uses domain-aware parameterization, ensuring that problems are not just mathematically solvable but are also grounded in physical and engineering realism.
-This is done by constraining all generated values by the principles of the domain being tested.
-For instance, in Chemical Engineering, problems use real reactants like Benzene, pre-validated stoichiometric equations, and temperature dependent heat capacity parameters rather than random constant values.
-In Electrical Engineering, parameters are grounded in fundamental constants (e.g., $c_{0}$, $\epsilon_{0}$) and real-world materials (e.g., Polyethylene), with signal properties procedurally generated to explicitly test concepts like aliasing.
-In Mechanical Engineering, problems co-select consistent properties for real materials (e.g., 6061-T6 Aluminum) and are context-aware, such as using Mercury for manometer problems.
-A comprehensive list of these parameterization examples is provided in Appendix E, and the authoritative data sources used are listed in Appendix F.
-Furthermore, across all templates, these already constrained parameters are subjected to multi-layer sanity checks.
-For example, a chemical engineering template verifies that a final concentration is less than the initial one.
-A mechanical engineering problem ensures a calculated stress does not exceed the material’s yield strength.
-Similarly, an electrical engineering template confirms a wave’s propagation speed does not exceed the speed of light.
-These checks validate that the final combination of values results in a solvable and physically meaningful scenario.
-4.5 AI-Assisted Quality Assurance
-To ensure the correctness and scalability of the benchmark, we developed an AI-Assisted Quality Assurance workflow.
-While traditional unit tests are effective, they proved impractical for our generative framework due to the high number of templates and the complexity of verifying physically plausible outputs.
-Our quality assurance workflow automates this process by employing an LLM as a first-pass Peer Reviewer to validate new problem templates before their inclusion in the final benchmark.
-The core of this workflow is a programmatic validation script that prepares a comprehensive validation package for the LLM.
-For each template, this package includes not only the template’s full Python source code but also three unique instances (question-solution pairs) generated by it.
-This context is crucial, as it allows the LLM to assess both the static logic of the code and the dynamic quality of its output.
-We then pass this package to the LLM via a carefully engineered prompt.
-As detailed in Appendix G, the prompt instructs the LLM to assume the persona of an expert engineering professor and to evaluate the template against a multi-axis rubric provided in Appendix H.
-Upon receiving the results, our system programmatically parses and evaluates the scores.
-A template, $T$, is automatically approved if and only if the score assigned to it for every core category ($c_{i}$) in our rubric meets or exceeds a predefined threshold, $\theta$.
-We formalize this approval condition as:
-$\forall c_{i}\in C,\text{Score}(T,c_{i})\geq\theta$
-We established a high threshold of $\theta=4$ based on empirical testing against a control set of known-good and deliberately flawed templates (Examples are provided in Appendix L).
-Any template failing to meet this criterion is automatically flagged for human review.
-This threshold demonstrated an optimal balance between maximizing automation and reliably catching potential errors.
-In practice, this automated review process proved highly effective.
-Out of the 90 total templates, only 5 (5.55%) were flagged by the LLM and required subsequent human review.
-[^1] This human review was performed by the authors, who possess relevant engineering domain expertise.
-While we did not calculate formal inter-rater reliability between the LLM and human reviewers for this quality assurance step, we manually inspected and corrected the flagged templates.
-4.6 Symbolic Generation and Combinatorial Diversity
-A core feature of EngChain is its programmatic generation, designed to create a problem space so vast that it effectively prevents evaluation through memorization.
-Rather than relying on a static set of problems, EngChain is built upon symbolic templates where parameters are systematically randomized within domain-aware, physical constraints.
-This design provides strong resistance to training data contamination, as it makes it statistically improbable that an LLM has encountered any specific problem instance.
-Consequently, successful performance on EngChain is stronger evidence of a model’s ability to generalize its reasoning capabilities.
-The model is forced to solve each problem from first principles rather than relying on pattern matching or rote recall of seen solutions.
+The core of EngTrace is the symbolic template, a Python blueprint for generating unique, physically grounded engineering problems.
+As shown in Figure 3, the pipeline first samples domain-aware parameters from authoritative data sources (Appendix E) and fundamental principles (see detailed examples in Appendix F).
+It then performs core computations with multi-layer sanity checks to derive computationally verified ground-truth traces.
+These are serialized into natural language (question, solution) pairs, creating a dynamic problem space that effectively resists data contamination.
+3.3 Template Validation
+To establish EngTrace as a high-fidelity “Gold Standard” benchmark, we utilize a robust two-stage validation process that pairs automated heuristic auditing with expert manual verification.
+Stage 1: Automated AI Pre-Screening.
+Building upon the “Panel of LLM” (PoLL) methodology (Verga et al., 2024), we construct an “AI Tribunal” comprising three frontier reasoning models: GPT-5 (OpenAI, 2025a), Claude Opus 4.5 (Anthropic, 2025b), and Gemini 3 (Google, 2025).
+Each model acts as an independent judge, utilizing the specific prompt provided in Appendix G, to evaluate the source code and generated instances against the strict multi-axis rubric detailed in Appendix H.
+To mitigate potential hallucinations and ensure output reliability, we apply a hybrid aggregation strategy requiring each template $T$ to satisfy three consensus conditions.
+First, the Median Quality Score ($S_{med}$) across the tribunal must reach a threshold $\theta=4$ for every evaluative dimension $d\in D$:
+$\forall d\in D,\quad S_{med}^{(d)}=\text{median}(S_{m_{1}}^{(d)},S_{m_{2}}^{(d)},S_{m_{3}}^{(d)})\geq 4$ (1)
+Second, we enforce a Majority Vote safety check, discarding any template where more than one judge identifies a violation ($\sum F_{i}\geq 2$, where $F\in\{0,1\}$).
+Finally, we quantify uncertainty using a Disagreement Score ($\sigma_{max}$), defined as the maximum standard deviation across all dimensions.
+Templates exhibiting high variance are flagged for manual review to resolve potential logical ambiguities:
+$\sigma_{max}=\max_{d\in D}\left(\sqrt{\frac{1}{3}\sum_{i=1}^{3}(S_{m_{i}}^{(d)}-\mu^{(d)})^{2}}\right)\leq 0.5$ (2)
+where $\mu^{(d)}$ represents the mean score across judges for dimension $d$.
+Stage 2: Expert Human Certification.
+Templates passing the AI tribunal undergo a double-blind review by domain experts through a custom Streamlit interface presented in Appendix I.
+We assess Inter-Annotator Agreement per branch using Fleiss’ Kappa ($\kappa$) for binary admissibility (Approve/Reject) and Krippendorff’s Alpha ($\alpha$) for ordinal quality ratings (1–5), the latter accounting for disagreement magnitude and missing data.
+AI-Human Alignment.
+To assess validation reliability, we measure correspondence between the AI Tribunal and human experts using False Positive Rate (FPR), to quantify filter safety (templates approved by AI but rejected by human experts), and Spearman’s Rank Correlation ($\rho$) to evaluate monotonic ranking consistency, chosen over Pearson to better handle the ordinal nature and high-quality skew of the data.
+Detailed results are provided in Appendix J.
+Figure 3: The EngTrace Template Generation Pipeline maps domain-aware constraints to serialized question-solution pairs via physically-validated computation, illustrated here using a CSTR volume calculation example.
 3.4 Difficulty Scaling
-Problem difficulty in EngChain is not a subjective label but a calibrated measure determined by systematically varying three primary factors:
-  • Conceptual Complexity refers to the number and integration of engineering principles required.
-An Easy problem typically tests a single, isolated concept like Ohm’s Law.
-An Intermediate problem requires the combination of two or three concepts, such as applying the principle of superposition to calculate a net electric field from multiple point charges.
-An Advanced problem demands the integration of multiple complex ideas, such as using numerical integration on Levenspiel plot data to determine a reactor volume.
-  • Mathematical Sophistication describes the nature of the required calculations.
-Easy problems involve direct algebraic substitution.
-Intermediate problems may require basic calculus or vector addition.
-Advanced problems escalate to solving differential equations or applying numerical methods, such as the trapezoidal rule for integration.
-  • Procedural Depth measures the length and interdependence of the reasoning chain.
-An Easy problem might be solvable in one or two steps.
-In contrast, an Advanced problem involves a long sequence of calculations where the output of each step is a critical input for the next, testing a model’s ability to maintain a coherent and accurate computational trace.
-This multi-axis framework for difficulty allows EngChain to move beyond a simple easy/hard dichotomy and provides a more fine-grained analysis of an LLM’s reasoning capabilities.
+Problem difficulty in EngTrace is calibrated through a multi-axis framework defined by three dimensions: Conceptual Complexity (isolated principles vs.
+multi-domain synthesis), Mathematical Sophistication (direct algebraic substitution vs.
+differential equations), and Procedural Depth (length of interdependent reasoning chains).
+A detailed quantitative breakdown of the dataset’s difficulty distribution is provided in Appendix K.
 4 Evaluation Framework
-For a verifiable assessment of the entire reasoning process, we compute several key metrics by comparing a model’s generated solution against the ground-truth reasoning trace.
-The evaluation is multifaceted.
-It jointly evaluates final answer accuracy, the fluency and semantic quality of the generated solution, and the correctness of the multi-step reasoning chain.
+Inspired by recent advances in process supervision (Lightman et al., 2023) and cascaded model verification (Chen et al., 2023), we introduce a two-stage evaluation framework that validates intermediate reasoning steps in addition to final-answer accuracy through a tiered verification protocol.
+To ensure alignment with professional standards, we validated this protocol through a blind human-expert study, confirming that our Tribunal-based approach significantly outperforms standard metrics in capturing engineering reasoning quality (see Appendix M for full methodology and ablation results).
+4.1 Preliminaries
+We define the evaluation of a reasoning chain as a comparison between the ground truth solution $G$ and the model-generated predicted solution $P$.
+Both are represented as ordered sequences of $M$ and $N$ discrete reasoning steps, respectively:
+$G=(g_{1},g_{2},\dots,g_{M}),\quad P=(p_{1},p_{2},\dots,p_{N})$ (3)
+Each individual reasoning step $s_{i}\in\{G,P\}$ consists of a textual derivation $t_{i}$ and, where applicable, an associated numerical or symbolic intermediate result $a_{i}=\text{Res}(s_{i})$.
+To assess the validity of the derivation, we seek a mapping that aligns steps in $P$ to their logical counterparts in $G$.
+Unlike standard string matching, this alignment must account for structural variations such as step merging, reordering, or alternative valid methodologies while ensuring that the cumulative logic remains faithful to the underlying engineering principles.
 4.2 Final Answer Accuracy
-It is determined by comparing the model’s predicted numerical answer ($A_{pred}$) to the ground-truth answer ($A_{gt}$).
-A response is considered correct only if its relative error is less than a strict tolerance ($\epsilon_{final}$) of 1%, a standard appropriate for engineering precision.
-The calculation is formalized as follows:
-$\text{Relative Error}=\left|\frac{A_{gt}-A_{pred}}{A_{gt}}\right|<\epsilon_{final}$
+To evaluate the correctness of the final predicted outcome, let $g_{M}$ and $p_{N}$ denote the last steps of the ground truth and predicted solutions, from which we extract the terminal numerical results $A_{G}=\text{Res}(g_{M})$ and $A_{P}=\text{Res}(p_{N})$.
+A response is deemed correct if the relative error between these values satisfies the following condition:
+$\mathbb{I}_{final}=\mathbb{I}\left(\frac{|A_{G}-A_{P}|}{|A_{G}|}<\epsilon_{final}\right)$ (4)
+where $\epsilon_{final}=0.02$ (2%), a strict threshold required to prevent catastrophic error propagation in high-precision engineering derivations.
 4.3 Textual Quality
-We assess the textual quality of a generated solution, evaluating its fluency and semantic fidelity against the ground-truth text using a combination of semantic and lexical similarity scores using the most common reference based evaluation metrics (Xie et al., 2023b):
-  • BERTScore (Zhang et al., 2020) measures deep semantic similarity by aligning contextual embeddings of tokens from the generated solution and the ground-truth solution.
-It computes precision, recall, and an F1-score to provide a robust measure of semantic equivalence, capturing nuances that lexical overlap metrics may miss.
-  • ROUGE-2 (Lin, 2004) quantifies the lexical overlap by calculating the F1-score based on the number of matching adjacent word pairs (bigrams) between the generated and ground-truth texts.
-Together, these scores provide a complementary assessment that reflects the model’s ability to produce solutions using appropriate and correctly structured domain-specific terminology.
-5.3 Multi-Step Reasoning Verification
-This is the core of our evaluation framework which assesses the entire logical chain of a model’s solution.
-This process moves beyond final answer accuracy by comparing each predicted reasoning step ($P_{j}$) against each ground-truth step ($GT_{i}$).
-A step is considered valid only if it satisfies two criteria: semantic similarity and numerical correctness.
-To formalize this, we compute a combined similarity matrix, $S_{comb}$, by taking the element-wise product (Hadamard product $\circ$) of a semantic similarity matrix ($S_{sem}$) and a numerical correctness matrix ($S_{num}$).
-$S_{comb}=S_{sem}\circ S_{num}$
-  • Semantic Similarity ($S_{sem}$) is computed using a powerful Cross-Encoder model (cross-encoder/stsb-roberta-large).
-This model processes each ground-truth and predicted step as a single contextual pair, yielding a high-fidelity similarity score that captures nuanced semantic relationships more effectively than bi-encoder models like the initially used SentenceTransformer.
-  • Numerical Correctness ($S_{num}$) is a binary matrix where an entry is 1 if the intermediate numerical answer in a predicted step matches the corresponding ground-truth step within a 2% relative error tolerance, and 0 otherwise.
-From this combined matrix, we determine if a step is “matched” by checking if its maximum similarity score exceeds a predefined threshold, $\tau$.
-We then calculate recall and precision to evaluate the reasoning chain.
-Recall measures the fraction of ground-truth steps the model successfully reproduced:
-$\text{Recall}=\frac{1}{|GT|}\sum_{i=1}^{|GT|}\mathbb{I}\left(\max_{j}S_{comb}[i,j]>\tau\right)$
-Precision measures the fraction of the model’s generated steps that were valid and relevant:
-$\text{Precision}=\frac{1}{|P|}\sum_{j=1}^{|P|}\mathbb{I}\left(\max_{i}S_{comb}[i,j]>\tau\right)$
-Finally, our primary metric for reasoning quality, the Step F1-Score, is calculated as the harmonic mean of the step-level recall and precision defined above, providing a single, robust measure of procedural correctness.
-$\text{Step F1-Score}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
-5.4 Automated Error Analysis with LLM-as-a-Judge
-While the Step F1-Score provides a sound quantitative measure of a model’s procedural correctness, it has a significant limitation: it can identify that a reasoning chain is flawed, but it cannot explain why.
-For instance, as will be discussed later, our preliminary results revealed models that achieved high final-answer accuracy despite very low reasoning scores, a classic case of being “right for the wrong reasons.” This made us realize that to diagnose these failures and determine their cause, a deeper, more qualitative analysis was required.
-Hence, to move beyond quantitative scores and introduce a layer of explainability, we propose a novel Automated Error Analysis system that employs a large language model as an expert evaluator, or an LLM-as-a-Judge.
-This system is triggered for any ground-truth step ($GT_{i}$) that fails to find a satisfactory match among the predicted steps.
-A step is flagged as a “failure” if its maximum combined similarity score falls below an empirically determined threshold, $\tau=0.7$.
-This condition is formalized as:
-$\max_{j}S_{comb}[i,j]<\tau$
-Figure 5: Automated error analysis with the LLM-as-a-Judge workflow.
-The overall workflow of this automated analysis process is illustrated in Figure 5.
-At its core, the process is rooted in carefully designed prompt engineering, where the judge model is given a system prompt defining its persona as an expert engineering professor and a set of critical rules for logical consistency.
-The input to the judge is a structured prompt containing the full problem context, the ground-truth step, and the model’s incorrect step.
-Providing this complete context is crucial for the judge to accurately classify the failure into one of the five distinct categories, which are defined in Appendix I.
-To ensure a reliably parsable output, the judge is instructed to return only a structured JSON object containing its classification and a detailed and scalable diagnosis of the model’s specific failure modes.
-The full system prompt used for the automated error analysis is provided in Appendix J.
+We evaluate the fluency and semantic fidelity of generated solutions using standard reference-based metrics (Xie et al., 2023b).
+We employ BERTScore (Zhang et al., 2020) to measure deep semantic similarity via contextual embedding alignment, and ROUGE-2 (Lin, 2004) to quantify lexical overlap.
+Together, these metrics assess the model’s ability to employ appropriate domain-specific terminology.
+4.4 Tiered Reasoning Verification
+To evaluate the full reasoning chain, we characterize the validity of the predicted sequence $P$ against the ground truth $G$ through a tiered verification protocol.
+This process consists of two phases: determining step-level validity through hierarchical assessment and aggregating these results into a global reasoning metric.
+Step-Level Validity Assessment.
+We compute a validity matrix $V\in\mathbb{R}^{M\times N}$, where each element $V_{i,j}$ represents the validity of predicted step $p_{j}$ relative to ground truth step $g_{i}$.
+This validity is determined by two successive tiers of verification.
+Tier 1: Automated Symbolic Verification.
+This tier fills the matrix $V$ by enforcing syntactic and numerical precision through a binary step-validity score:
+$S_{\mathcal{T}_{1}}(g_{i},p_{j})=\delta_{num}\land\delta_{sem}$ (5)
+where $S_{\mathcal{T}_{1}}=1$ only if the numerical consistency (relative error tolerance $\epsilon=0.02$) and semantic similarity (Cross-Encoder threshold $\tau=0.7$) are both satisfied.
+To optimize computational efficiency, we calculate the alignment ratio, which represents the proportion of steps in $P$ successfully verified by Tier 1.
+If this ratio exceeds a confidence threshold ($\rho=0.8$) and the final answer is correct, the reasoning chain is classified as a standard derivation and accepted without further review.
+Chains falling below this threshold are escalated to Tier 2 for heuristic assessment.
+Tier 2: Multi-Model Heuristic Verification.
+Escalated chains are evaluated by a heterogeneous AI Tribunal to identify valid alternative derivations (prompt provided in Appendix L).
+Each judge $J_{k}$ classifies the observed discrepancy between $p_{j}$ and $g_{i}$ into a category $C_{k}$, which is then mapped to a scalar score:
+$\mu(C)=\begin{cases}1.0&\text{if }C=\text{Alternate Correct}\\ 0.5&\text{if }C=\text{Calculation Error}\\ 0.0&\text{if }C\in\{\text{Conceptual Error, Other}\}\end{cases}$ (6)
+The majority consensus of these scalar scores defines $S_{\mathcal{T}_{2}}$, which is assigned to $V_{i,j}$ to finalize the validity matrix for non-standard derivations.
+In the absence of a clear majority, the assessment defaults to a conservative tie-breaker where the lowest score is assigned.
+Global Reasoning Metric $\mathbf{F1_{rec}}$.
+To quantify the overall quality of the reasoning chain, we calculate the Recovered Reasoning F1 Score ($F1_{rec}$).
+We first identify the optimal alignment that maximizes the aggregate verification score:
+$\mathcal{R}_{total}=\max_{\sigma\in\mathcal{A}}\sum_{(i,j)\in\sigma}V_{i,j}$ (7)
+where $\mathcal{A}$ represents the set of valid one-to-one mappings between steps.
+This optimization is solved using the Hungarian matching algorithm on the finalized validity matrix $V$.
+We define recovered precision and recall as $P_{rec}=\mathcal{R}_{total}/N$ and $R_{rec}=\mathcal{R}_{total}/M$, respectively, where $N$ and $M$ are the number of predicted and ground truth steps.
+Finally, $F1_{rec}$ is computed as the harmonic mean of $P_{rec}$ and $R_{rec}$.
 5 Experiments
 5.1 Evaluation Model Suite
-To provide a comprehensive evaluation of engineering reasoning capabilities, we evaluate models from a diverse suite.
-The results presented in Section 7 focus on 11 distinct models from the first category.
-While a full evaluation of the other categories is planned for future work, we group the complete suite of models into the following four-way taxonomy:
-  1. 1.
-    Frontier proprietary models, which represent the current upper bound of performance.
-This includes models such as GPT-{5, 4.1, 5-mini, 4.1-mini} (OpenAI, 2025a, b), Claude Sonnet {4.5, 4, 3.7} (Anthropic, 2025b, c, a), Gemini 2.5 {Pro, Flash} (Comanici et al., 2025), DeepSeek {V3.2, V3.1, R1} (Liu et al., 2024; Guo et al., 2025), and Grok 4 {Heavy, Fast} (xAI, 2025).
-  2. 2.
-    Science/Engineering-enhanced models, such as Galactica (Taylor et al., 2022), Minerva (Lewkowycz et al., 2022), ChemLLM (Zhang et al., 2024), SciLit01 (Li et al., 2025a), and SciAgent (Gao et al., 2024), are systems pre-trained or fine-tuned on vast scientific and technical corpora to improve domain-specific knowledge and complex procedural reasoning.
-  3. 3.
-    Math-enhanced models, such as WizardMath (Luo et al., 2023), MetaMath (Yu et al., 2023), Mathstral (Mistral, 2024), and Qwen2.5-Math (Yang et al., 2024), are systems specifically trained on mathematical and symbolic corpora to enhance their quantitative reasoning capabilities.
-  4. 4.
-    General-purpose open models, such as LLaMA 3.1 (Grattafiori et al., 2024) and Qwen {2.5, 3} (Qwen, 2024, 2025), provide strong, domain-agnostic foundations that are often used to create specialized variants.
-This four-way taxonomy provides a structure for comparing different model architectures and training paradigms, highlighting how model scale, domain specialization, and mathematical supervision influence symbolic engineering reasoning.
-The detailed configurations and model sources can be found in Appendix K.
+We evaluate 24 models categorized into a three-way taxonomy to isolate the impacts of scale, access, and specialization: (1) Frontier proprietary models represent the current state-of-the-art in reasoning and instruction following.
+Our evaluation includes 12 models in this category: GPT-5 {Base, Mini} and GPT-4.1 {Base, Mini} (OpenAI, 2025a, b), Claude Sonnet {4.5, 4, 3.7} (Anthropic, 2025c, d, a), Gemini {3 Pro, 2.5 Pro, 2.5 Flash} (Google, 2025; Comanici et al., 2025), and DeepSeek {V3, R1} (Liu et al., 2024; Guo et al., 2025a).
+(2) General-purpose open models provide strong, domain-agnostic foundations and allow for transparent analysis of performance relative to parameter count.
+We evaluate 8 models: Meta-Llama-3.1 {8B, 70B} (Grattafiori et al., 2024), Qwen2.5 {7B, 14B, 72B} and Qwen3-8B (Qwen, 2024, 2025), and Gemma {2-9B, 3-27B} (Gemma Team et al., 2024; Gemma Team and Google DeepMind, 2025).
+(3) Math-enhanced models are systems specifically trained on mathematical and symbolic corpora to enhance quantitative reasoning.
+We evaluate 4 specialized models in the 7B parameter class: Qwen2.5-Math-7B (Yang et al., 2024), Mathstral-7B (Mistral, 2024), WizardMath-7B (Luo et al., 2023), and MetaMath-7B (Yu et al., 2023).
+Detailed configurations and model sources can be found in Appendix N.
 5.2 Experimental Setup
 We instantiate the benchmark by sampling 15 instances per symbolic template, each with a distinct random seed, creating a total of 1,350 test cases (90 templates $\times$ 15 instances) for each model.
-We evaluate all models under a unified decoding configuration: a low temperature of 0.2 and a maximum token limit of 4,096.
-We use a low temperature (T=0.2) to ensure deterministic, high-confidence outputs and a high token limit (4,096) to allow ample space for detailed, multi-step solutions without premature cut-offs.
-We use a zero-shot setup with the following standardized reasoning prompt, which is used for all inference tasks:
-We use a zero-shot setup with the following standardized reasoning prompt, which is used for all inference tasks:
-You are an expert engineer. Solve the following
-problem by providing a detailed, structured
-solution. Use the exact headings and
-formatting provided below.
-# Given
-List all known variables and their
-values with units.
-# Find
-State the variable(s) to be calculated.
-# Formulae
-Write down all necessary governing equations
-before substituting any values.
-# Solution
-Provide a step-by-step calculation. Each step
-must start on a new line and be formatted exactly
-as ’**Step X:**’, where X is the step number.
-Show the substitution of values into
-the formulae clearly.
-# Final Answer
-State the final numerical result with its units
-in the format: **Answer:** [value] [units]
-To process the model outputs, we first segment the text into reasoning steps.
-We use a regular expression to identify line-start markers that match either the Step <number> format (e.g., **Step 1**) or the numbered-list format (e.g., 1.).
-A subsequent regex cleans these prefixes.
-We then extract numerical values from each step and the final answer using a three-priority heuristic: first, matching the **Answer:** tag; second, falling back to the last number following an equals sign (e.g., V = 100.0); and third, taking the last standalone number.
-If no overall **Answer:** tag is found, the final numerical value from the last step is used for verification.
-7 Results
-7.1 Overall Model Performance
-The primary finding from our evaluation is a stark and universal divergence between the models’ ability to produce a correct final answer and the verifiable correctness of their reasoning process.
-As shown in  Table 1, while the top models achieve a final answer accuracy of approximately 63.1% (gpt-4.1-mini and gpt-5-mini), their procedural reasoning capability is critically low.
-The Reasoning (F1 Score), our primary metric for procedural correctness, is alarmingly poor across the board, with the top-performing model, gemini-2.5-flash, only achieving 19.32%.
-This demonstrates that in the vast majority of cases, all models fail to follow a sound, verifiable problem-solving methodology, confirming the “right for the wrong reasons” phenomenon at scale.
-This significant reasoning gap is further masked by high textual fluency.
-The BERTScore, for instance, remained consistently high for all models, averaging approximately 0.87.
-This indicates that models excel at generating semantically appropriate, engineering-domain language, even when the underlying procedural logic is flawed or nonsensical.
-Model Final Answer Accuracy $\uparrow$ Reasoning (F1 Score) $\uparrow$ BERTScore $\uparrow$ ROUGE-2 $\uparrow$ ROUGE-L $\uparrow$
-Frontier Proprietary Models
-Claude 3.7 56.8149.55 16.9121.76 88.031.16 29.287.65 37.097.63
-Claude 4 57.7849.41 14.1618.28 87.931.33 29.558.16 37.257.91
-Claude 4.5 52.2249.97 14.1919.72 87.391.09 28.108.15 35.347.58
-Deepseek v3.2 61.2648.73 13.5319.54 86.541.23 25.698.96 34.579.12
-Deepseek R1 59.9349.02 19.0324.82 86.481.34 25.588.34 34.498.30
-Gemini 2.5 Flash 59.4149.13 19.3223.56 87.551.63 28.829.17 36.409.36
-Gemini 2.5 Pro 59.9349.02 17.9721.01 88.241.41 31.828.29 37.918.49
-GPT 4.1 57.7849.41 12.1315.85 86.651.24 26.749.13 34.459.12
-GPT 4.1 Mini 63.1148.27 14.8419.69 86.521.14 25.198.17 33.828.34
-GPT 5 60.4448.92 15.4721.79 88.191.21 28.988.30 36.558.06
-GPT 5 Mini 63.1148.27 14.3020.32 87.951.10 26.637.62 34.747.52
-Table 1: Overall performance of frontier models on the EngChain benchmark (N=1350) under a Zero-Shot setting.
-Scores represent the mean, with standard deviation shown in superscript.
-The highest mean score in each column is in bold.
-7.2 Performance by Engineering Branch
-A fine-grained analysis of the results reveals a significant performance disparity across the three engineering branches, as illustrated in Figure 6.
-We observe that model reasoning capabilities are not uniform and are highly sensitive to the specific scientific principles and mathematical structures of a given engineering domain.
-Chemical Engineering consistently proved to be the most challenging domain for all models, yielding the lowest scores in both final answer accuracy and procedural reasoning.
-The reasoning collapse in this domain is particularly severe, with several models falling into single-digit Reasoning (F1 Score)s, including Deepseek v3.2 (6.90%) and GPT 5 Mini (7.58%).
-Conversely, Mechanical Engineering was clearly the most solvable domain, where models achieved their highest scores.
-This branch produced the peak Final Answer Accuracy of the entire benchmark (Gemini 2.5 Pro at 70.22%) and the highest Reasoning (F1 Score) (Deepseek R1 at 30.11%).
-Electrical Engineering demonstrated intermediate performance, with scores that consistently fell between the other two branches.
-Its peak Reasoning (F1 Score) of 22.49% (Gemini 2.5 Flash) and peak Final Answer Accuracy of 67.11% (GPT 4.1 Mini) were modest, clearly lagging behind the top scores in Mechanical Engineering but remaining well above the challenging baseline set by Chemical Engineering.
-Figure 6: Branch-level performance comparison of the top four models ranked by mean Reasoning (F1 Score) across the three engineering branches.
-Chemical Engineering is consistently the most challenging domain for procedural reasoning.
-7.3 Performance by Engineering Domain
-Further analysis reveals that model performance is highly inconsistent across the nine distinct engineering domains, as visualized in the radar plot in Figure 7.
-This “spiky” performance profile, where a model can excel in one domain while failing in another, suggests that current models possess patches of specialized knowledge rather than generalized, first-principles reasoning.
-The data shows several domains are exceptionally challenging for all models.
-Domains such as Reaction Kinetics, Mechanics of Materials, Transport Phenomena, and Electromagnetics & Waves produced critically low reasoning scores.
-In these areas, it was common for top-tier models to score in the single digits, with Deepseek v3.2 scoring 6.90% in Reaction Kinetics and GPT 5 Mini scoring 7.58% in Mechanics of Materials.
-Conversely, the domain of Vibrations & Acoustics emerged as a clear outlier, representing the easiest domain for nearly all models.
-The reasoning scores here were dramatically higher than in any other domain, with Deepseek R1 achieving 52.26% and Gemini 2.5 Flash achieving 40.70%.
-The domain of Signals & Systems was another area of relative strength, where most models achieved some of their highest reasoning scores.
-This domain-level variance underscores the need for broad, multi-domain benchmarks.
-We hypothesize this performance gap is linked to the underlying mathematical complexity.
-The high-scoring domains in our results often feature problems solvable with direct algebraic substitution or linear, discrete procedures.
-The most challenging domains, in contrast, frequently require models to set up and solve differential equations or manage systems of complex, interdependent variables, which appear to be critical failure points for current procedural reasoning.
-Figure 7: Domain-level performance comparison of the top four models ranked by mean Reasoning (F1 Score).
-The plot reveals a “spiky” performance profile, highlighting that model reasoning is highly specialized and not generalized across topics.
-7.4 Performance by Problem Difficulty
-As expected, model performance degrades as problem complexity increases, but this trend is most pronounced in Final Answer Accuracy.
-Most models maintain a respectable accuracy of 60-70% on Easy and Intermediate tasks.
-However, performance drops sharply on Advanced problems, with nearly all models falling into the 40-50% range.
-For instance, gemini-2.5-pro’s accuracy falls from 68.6% on Intermediate problems to just 45.2% on Advanced.
-The Reasoning (F1 Score), by contrast, tells a more complex story, as illustrated in Figure 8.
-Instead of a uniform decline, the reasoning scores for many top models (such as claude-3-7 and gpt-5) are lowest on Intermediate tasks before recovering slightly on Advanced tasks.
-A few models (e.g., gemini-2.5-flash and deepseek-reasoner) even show a slight improvement in reasoning as difficulty increases.
-The most critical finding, however, is that reasoning scores remain universally low (sub-21%) across all three difficulty levels.
-This suggests the models’ procedural reasoning failures are fundamental and not simply a function of task complexity.
-Figure 8: Difficulty-level performance comparison of the top four models ranked by mean Reasoning (F1 Score).
-The plot shows procedural reasoning performance is critically low across all levels of complexity, not just on advanced tasks.
-7.5 Qualitative Error Analysis with LLM-as-Judge
-While the quantitative Reasoning (F1 Score) identifies that a reasoning chain is flawed, it cannot explain why.
-To diagnose the root causes of these failures, we employed our LLM-as-Judge system (described in Section 5) to perform a qualitative error analysis on Gemini 2.5 Pro.
-We selected this model as our representative subject because it is a strong “all-rounder,” achieving both a top-tier reasoning score (17.97%) and the first rank on all language fluency metrics (BERTScore, ROUGE-2, and ROUGE-L).
-The analysis was done using GPT-4o (OpenAI, 2024) as the judge or expert evaluator model.
-The judge was used to analyze all 4,118 reasoning steps that were flagged by our verifier as a “mismatch” from the ground-truth solution.
-The results, summarized in Table 2, are striking.
-The single most significant finding is that the vast majority of flagged steps were not errors at all.
-Our judge classified 73.94% (3,045 instances) as Alternative Correct or in other words, valid reasoning paths that simply differed from the ground-truth solution.
-This overwhelmingly indicates that models are frequently penalized by rigid, single-path evaluations for finding a valid, creative, or different solution, a phenomenon masked by low quantitative scores.
-For the remaining 1,072 instances identified as genuine errors, Conceptual Error was the dominant failure mode, accounting for 59.1% of all true errors (634 instances).
-This demonstrates that models are far more likely to fail because they misapply a physical principle or choose the wrong formula, rather than making simple Calculation Errors (14.3%) or Input Errors (7.6%).
-This finding pinpoints the primary weakness of modern LLMs not in their arithmetic, but in their application of deep, domain-specific knowledge.
-A full breakdown of error distributions by branch, domain, and difficulty can be found in Appendix M.
-Error Category Total Count Percentage (%)
-Alternative Correct 3,045 73.94%
-Conceptual Error 634 15.40%
-Other 204 4.95%
-Calculation Error 153 3.72%
-Input Error 81 1.97%
-Analysis Failed 1 0.02%
-Total 4,118 100.0%
-Table 2: Overall distribution of error types identified by Gemini 2.5 Pro as LLM-as-Judge across all 4,118 flagged reasoning mismatches.
+We evaluate all models under a unified decoding configuration: a low temperature of 0.2 to ensure deterministic outputs and a maximum token limit of 4,096 to allow for detailed reasoning.
+All models are evaluated in a zero-shot setting using a standardized prompt and output parsing pipeline; full implementation specifications are provided in Appendix O.
+5.3 Results
+5.3.1 Overall Model Performance
+Table 1 summarizes performance across five key metrics: Final Answer Accuracy (FAC), Reasoning F1, and standard semantic measures (BERTScore, ROUGE-2/L).
+The results reveal a distinct hierarchy: frontier proprietary models define the state-of-the-art, with Gemini 3 Pro achieving peak numerical precision (FAC 64.00%) and DeepSeek R1 leading in logical process alignment (Reasoning F1 44.41%).
+In contrast, open-weights models lag significantly in both metrics despite competitive semantic scores, even among large-scale systems like Llama 3.1 70B.
+Notably, specialized math-enhanced models in the 7B parameter class generally underperform general-purpose counterparts of equivalent size, suggesting that abstract mathematical pre-training does not strictly transfer to the physically constrained reasoning required for engineering.
+Model Final Answer $\uparrow$ Reasoning (F1) $\uparrow$ BERTScore $\uparrow$ ROUGE-2 $\uparrow$ ROUGE-L $\uparrow$
+Frontier Proprietary LLMs
+Claude 3.7 Sonnet 57.63 39.36 88.09 29.42 37.36
+Claude 4 Sonnet 56.81 36.80 87.92 29.52 37.22
+Claude 4.5 Sonnet 51.78 33.61 87.37 28.06 35.39
+DeepSeek V3 62.37 37.06 86.54 25.69 34.57
+DeepSeek R1 61.04 44.41 86.48 25.58 34.49
+Gemini 2.5 Flash 59.78 42.22 87.64 29.14 36.70
+Gemini 2.5 Pro 60.44 39.38 88.24 31.88 38.07
+Gemini 3 Pro 64.00 37.05 86.71 27.32 34.99
+GPT-4.1 57.11 31.22 86.63 26.78 34.46
+GPT-4.1 Mini 61.56 36.87 86.50 25.06 33.87
+GPT-5 61.04 41.81 88.19 28.98 36.55
+GPT-5 Mini 63.70 41.01 87.95 26.63 34.74
+General Purpose Open LLMs
+Gemma 2 9B 26.59 21.37 88.19 28.02 35.77
+Gemma 3 27B 44.30 27.59 87.04 23.93 32.66
+Llama 3.1 8B 23.70 18.61 87.65 25.59 32.50
+Llama 3.1 70B 40.81 30.04 88.34 28.87 35.82
+Qwen 2.5 7B 35.63 23.67 86.30 24.21 32.89
+Qwen 2.5 14B 42.07 27.67 86.44 24.94 33.38
+Qwen 2.5 72B 48.07 28.02 86.46 25.58 33.96
+Qwen 3 8B 23.70 16.68 86.91 8.93 11.09
+Math Enhanced LLMs
+Mathstral 7B 19.70 18.59 87.79 28.02 36.42
+MetaMath 7B 6.07 2.37 82.97 10.57 16.69
+Qwen 2.5 Math 7B 23.04 18.18 85.61 21.71 29.98
+WizardMath 7B 8.22 6.38 83.66 14.14 20.38
+Table 1: Overall Performance Comparison of 24 models on the EngTrace benchmark.
+Scores are reported as percentages.
+The Final Answer column denotes strict numerical accuracy, while Reasoning (F1) measures the structural alignment of the derivation steps.
+The highest score in each column is highlighted in bold.
+5.3.2 Performance by Engineering Branch
+Analysis by branch reveals a trade-off between reasoning fidelity and numerical precision.
+Chemical Engineering proved the most rigorous, yielding the lowest average scores; DeepSeek R1 achieved a decisive lead here (FAC 64.89%, F1 40.91%), confirming that complex stoichiometry demands sustained chain-of-thought capabilities.
+In Electrical Engineering, Gemini 3 Pro secured the top FAC (68.67%), while DeepSeek R1 retained the highest Reasoning F1 (41.22%), indicating structurally superior derivations even when numeric outputs diverge.
+Mechanical Engineering was the most accessible, with Gemini 2.5 Pro achieving the benchmark’s highest overall FAC (71.56%).
+Notably, Llama 3.1 70B showed strong semantic performance here (BERTScore 88.37%), suggesting mechanical terminology is particularly well-represented in general pre-training corpora.
+5.3.3 Performance by Engineering Domain
+A granular analysis of specific domains (Figure 4) reveals critical variances in capability.
+In Chemical Engineering, DeepSeek R1 dominates Thermodynamics and Reaction Kinetics (FAC 78.33% and 58.00%), suggesting its reinforcement learning optimization effectively navigates coupled equations; conversely, Transport Phenomena remains a bottleneck (max 53.33% FAC).
+Electrical Engineering results are heterogeneous: Gemini 3 Pro excels in Digital Communications (77.33% FAC), likely due to its logic-based nature, but physics-heavy domains like Electromagnetics remain challenging, with no model exceeding 56.00% FAC.
+In Mechanical Engineering, Mechanics of Materials appears largely solved (Gemini 3 Pro: 92.67% FAC), while Fluid Mechanics caps at 56.00% FAC.
+This indicates that while models have mastered static relationships, they continue to struggle with the non-linear differential equations governing fluid dynamics.
+Figure 4: Domain-Level Performance comparison of the top four models ranked by mean Final Answer Accuracy.
+The plot reveals a “spiky” performance profile, highlighting that model capabilities are highly specialized and not generalized across topics.
+5.3.4 Performance by Difficulty Level
+Stratifying performance by difficulty (Figure 5) reveals a functional transition in model utility.
+In the Easy tier, efficiency-oriented architectures like GPT-5 Mini (67.84%) and open models like Llama 3.1 70B (52.55%) perform considerably well, suggesting flagship compute is unnecessary for fundamental engineering problems.
+However, the Advanced tier exposes a sharp divergence: Llama 3.1 70B collapses to 23.33%, identifying frontier systems as the sole viable option for high-complexity scenarios, led by Gemini 3 Pro (54.55%).
+Notably, specialized math-enhanced models (e.g., Qwen 2.5 Math) consistently underperform, peaking at just 32.35% on Easy tasks, confirming that abstract mathematical pre-training is insufficient for domain-specific physical reasoning.
+Figure 5: Difficulty-Level Performance of the top four models by mean Final Answer Accuracy.
+The plot shows a “complexity cliff”: efficiency-oriented models degrade sharply at advanced tasks, whereas frontier reasoning models remain stable.
+5.4 Error Analysis
+To characterize reasoning failures beyond simple accuracy metrics, we analyzed the 17,528 reasoning chains (54.1% of the dataset) that failed Tier 1 verification.
+This audit reveals that across all models, 85.86% of traces flagged as discrepancies were identified by the AI Tribunal as Alternative Correct derivations, as illustrated in Figure 6.
+This rate approached saturation in frontier models (97.17%), confirming that advanced models possess a high degree of solution diversity that renders rigid template-matching metrics unsuitable for engineering evaluation.
+Focusing on confirmed failures, the aggregate data shows that Conceptual Errors (1,318) outnumbered Calculation Errors (711) by a ratio of nearly 2:1.
+This indicates that while models frequently hallucinate physics principles, they are comparatively more reliable in arithmetic calculations.
+This trend correlates strongly with complexity: while Easy tasks exhibited a balanced error distribution, Advanced tasks saw the conceptual-to-calculation failure ratio widen to 3.7:1.
+Finally, 795 “Lucky Guesses” (2.5%), defined as correct answers with structurally invalid reasoning, underscore the necessity of multi-stage verification over final-answer accuracy alone.
+Figure 6: Distribution of Identified Reasoning Errors across model categories.
+While Frontier models primarily utilize valid alternative methods, General Open and Math-Enhanced models exhibit a significantly higher density of confirmed reasoning failures.
 6 Conclusion and Future Work
-We introduced EngChain, a symbolic benchmark for verifiable, multi-step engineering reasoning.
-Built from a taxonomy of 90 problem templates across 9 domains and 3 engineering branches, EngChain provides a verifiable evaluation of multi-step engineering problem-solving, assessing the entire reasoning chain rather than only the final answer.
-Our comprehensive evaluation of frontier models reveals a stark “right for the wrong reasons” phenomenon: a universal gap between modest Final Answer Accuracy and critically low Reasoning (F1 Score)s.
-Our qualitative analysis further reveals that this gap is not only due to error, but that a striking majority of flagged mismatches are actually Alternative Correct solutions, which single-path verifiers would wrongly penalize.
-For genuine failures, we find that Conceptual Errors are the dominant failure mode, not simple calculation mistakes.
-Future work will proceed in several key directions.
-First, we will complete our planned evaluation by benchmarking the science-enhanced, math-enhanced, and general-purpose open models to provide a holistic comparison across all model categories.
-Second, we will expand the LLM-as-Judge analysis to other top-performing models to create diagnostic profiles of their unique failure modes.
-Third, to rigorously validate our qualitative methodology, we will conduct two crucial follow-up studies.
-A Human Evaluation Study will be performed to establish a definitive ground-truth for our judge’s classifications.
-This will be complemented by a Cross-Judge Consistency study to ensure our error categorizations are not an artifact of our chosen evaluator (GPT-4o (OpenAI, 2024)).
-This consistency study will involve a diverse set of judges, including a model from a different leading developer, such as Claude 4 Opus (Anthropic, 2025c), and a high-capability open-source model, such as Qwen 2.5 (Qwen, 2024).
-Finally, upon completion of these studies, all evaluation scripts, prompts, and detailed results will be publicly released to ensure full reproducibility.
+In this work, we introduced EngTrace, a symbolic benchmark designed to enable verifiable process supervision of physically grounded engineering reasoning in LLMs.
+Our evaluation of 24 models reveals a “complexity cliff,” where efficiency-oriented and open-weights models degrade steeply in advanced engineering tasks while frontier reasoning models maintain stability.
+These findings highlight a critical reasoning gap, emphasizing that trustworthy engineering AI requires more than just general logic.
+Future work will explore tool-augmented reasoning to decouple semantic physical logic from computational precision and investigate uncertainty quantification to determine if models can reliably calibrate confidence in safety-critical engineering scenarios.
 Limitations
 We acknowledge several limitations in this work that we plan to address in future research.
-First, a significant limitation lies in our primary quantitative metric, the Reasoning (F1 Score).
-By design, this metric measures alignment against a single, pre-defined ground-truth path.
-However, our own qualitative analysis (Section 7.5) revealed that a striking majority of steps flagged as “mismatches” were actually Alternative Correct solutions.
-This indicates that our quantitative F1 score, when used in isolation, significantly underestimates the true reasoning capabilities of the models and inherently penalizes valid, alternative problem-solving approaches.
-Future work should therefore focus on developing more advanced evaluation methods, such as semantic alignment metrics or dynamic, model-based verifiers, which can programmatically accept and validate multiple correct reasoning paths.
-Second, EngChain is entirely synthetic, generated from symbolic templates.
+First, our current evaluation treats models as isolated inference engines, detached from the practical constraints of applied engineering.
+By focusing solely on static text generation, we overlook two critical components of reliable design: iterative self-correction and risk assessment.
+Consequently, while we effectively measure raw theoretical proficiency, our current framework does not fully capture the stability or safety assurance required for deployment in high-stakes engineering workflows.
+Second, EngTrace is entirely synthetic, generated from symbolic templates.
 While this design ensures verifiability and contamination resistance, it does not capture the linguistic diversity or contextual ambiguity of real-world engineering.
@@
 It does not yet include other major fields, such as civil, aerospace, or biomedical engineering.
-A clear next step is to expand the EngChain taxonomy in collaboration with domain experts to cover these additional fields, ensuring a more comprehensive test of engineering knowledge.
+A clear next step is to expand the EngTrace taxonomy in collaboration with domain experts to cover these additional fields, ensuring a more comprehensive test of engineering knowledge.
+Ethical Statement and Broad Impact
+This work utilizes synthetic data generated through symbolic templates and standard engineering principles; no private, sensitive, or proprietary industrial data was used.
+While EngTrace aims to advance reliable engineering reasoning, we emphasize that current LLMs should not be deployed in safety-critical physical systems without rigorous, hardware-in-the-loop verification.
+We hope this benchmark fosters the development of interpretable, verifiable, and physically grounded AI for high-stakes engineering applications.
+Data License
+The EngTrace dataset and associated code will be released under the MIT License.
+Acknowledgments
+We extend our gratitude to the engineering professionals and domain experts for validating the EngTrace framework and reviewing the templates.
 Appendix A LLM Prompt for Domain Validation
+To identify core domains, we consulted a panel of four frontier models: GPT-4o, Claude 3.5 Sonnet, Gemini 2.5 Pro, and DeepSeek V3.
+Each was prompted individually to identify cornerstone engineering domains.
+We finalized the nine core domains via majority voting.
+The domain identification prompt is provided below.
 You are an expert academic and senior curriculum designer at a top-tier engineering university (like MIT or Caltech).
@@
 } Do not include any preamble, conversational text, explanations, or markdown formatting around the JSON block.
-The above prompt was used to query large language models to cross-reference and validate the selection of fundamental engineering domains identified in our literature review.
 Appendix B LLM Prompt for Area Validation
-The following prompt was used to query large language models to identify the fundamental areas or sub-topics within each core engineering domain.
+Following the domain validation procedure, we utilized the same model panel and aggregated the pedagogical areas via majority voting. The prompt used for area identification is provided below.
 You are an expert academic and senior curriculum designer at a top-tier engineering university (like MIT or Caltech).
@@
 Do not include any preamble, conversational text, explanations, or markdown formatting around the JSON block.
-Appendix C Foundational Textbooks for Problem Selection
+Appendix C Foundational Textbooks for Template Selection
 The core engineering principles, problem typologies, and realistic constraints for the templates developed by us were sourced from the authoritative textbooks listed in Table 2.
-These texts, which are standard in top-tier engineering curricula, served as the primary knowledge base to ensure EngChain is pedagogically sound and grounded in the core engineering curriculum.
+These texts, which are standard in top-tier engineering curricula, served as the primary knowledge base to ensure EngTrace is pedagogically sound and grounded in the core engineering curriculum.
 They were also used to validate our selection of fundamental domains and areas.
@@
 Appendix D LLM Prompt for Pedagogical Scoring
-The following prompt was used to query large language models (specifically, gemini-2.0-flash) to evaluate the pedagogical significance of each sub-topic within its parent domain.
+The following prompt was used to query LLMs (described in Section A) to evaluate the pedagogical significance of each sub-topic within its parent domain.
 You are an expert academic and senior curriculum designer at a top-tier engineering university with decades of experience structuring undergraduate programs for ABET accreditation.
@@
 } Do not include any preamble, conversational text, explanations, or markdown formatting around the JSON block.
-Appendix E Detailed Domain-Aware Parameterization Examples
-To ensure that problems are not just mathematically solvable but are also grounded in physical and engineering realism, EngChain constrains all generated values by the principles of the domain being tested.
-The following provides specific examples of this methodology for each engineering branch.
-E.0.1 Chemical Engineering
-  • Reaction Kinetics and Stoichiometry: Instead of abstract chemicals, problems select from curated lists of common gas, liquid, and biochemical reactants like Propane, Benzene, and Glucose.
-More importantly, many problems are built around pre-validated, balanced chemical equations to ensure that the fundamental principles of mass conservation and stoichiometry are carefully upheld from the outset.
-  • Thermodynamics: Problems involving thermodynamic calculations use a multi-layered approach to realism.
-For common substances, a complete and consistent set of critical properties ($T_{c},P_{c},V_{c}$) is retrieved.
-For reaction calorimetry, problems utilize tabulated standard heats of formation ($\Delta H_{f}^{\circ}$).
-Crucially, for heat transfer calculations, the model moves beyond constant specific heats by incorporating temperature-dependent heat capacity parameters ($C_{p}/R=A+BT+CT^{2}+DT^{-2}$), ensuring a high degree of physical accuracy across a range of conditions.
-  • Transport Phenomena and Fluid Mechanics: To ensure realism in fluid dynamics problems, templates draw from extensive tables of fluid properties, including the density and viscosity of common liquids and gases.
-The benchmark also includes parameters for non-Newtonian fluids (e.g., Polymer Solutions & Melts etc), allowing it to test more complex, real-world rheological behaviors beyond simple Newtonian assumptions.
-E.0.2 Electrical Engineering
-  • Electromagnetics and Waves: Instead of using an arbitrary number, the phase velocity of a wave is determined by selecting a specific propagation medium from a curated list (e.g., Distilled Water, Polyethylene, Fused Silica etc).
-This ensures that all subsequent calculations for parameters like wavelength or wave number are physically consistent with that material’s properties.
-The generation process is further constrained by fundamental physical constants, including the speed of light in a vacuum ($c_{0}$) and the permittivity of free space ($\epsilon_{0}$).
-  • Signals and Systems: In scenarios bridging continuous and discrete-time domains, such as sampling, the sampling frequency is intentionally set well above the Nyquist rate to model ideal conversion and isolate the effects of the discrete-time system.
-For purely discrete-time problems, parameters are procedurally constrained to test specific concepts.
-For instance, signal frequencies are generated as rational multiples of $\pi$, and downsampling problems are deliberately constructed to either cause or avoid aliasing, directly testing the understanding of frequency folding.
-For fundamental operations like convolution, sequences are kept short and integer-valued to focus the analysis on the computational process.
-  • Digital Communications: For deterministic and random signal analysis, probabilities for discrete random variables are generated as exact rational numbers that sum perfectly to one.
-In digital modulation scheme problems, carrier frequencies are constrained to be integer multiples of the bit rate to simplify analysis, energy per bit ($E_{b}$) and signal-to-noise ratios ($E_{b}/N_{0}$) are generated in physically typical ranges (e.g., picojoules and realistic dB values), and modulation orders ($M$) are restricted to powers of two.
-Furthermore, data rates are assigned practical units such as kbps, Mbps, or Gbps, grounding the problems in contemporary engineering contexts.
-E.0.3 Mechanical Engineering
-  • Mechanics of Materials: Problems are grounded in the behavior of real-world materials by co-selecting properties from comprehensive tables.
-When a material such as “6061-T6 Aluminum” or “Carbon Fiber Reinforced Polymer” is chosen, its physically consistent Young’s Modulus ($E$), Poisson’s Ratio ($\nu$), and Shear Modulus ($G$) are used for all calculations.
-This ensures that problems involving stress, strain, and torsional deformation are realistic.
-Properties are available in both SI (GPa) and US Customary (ksi) units to reflect practical engineering work.
-  • Fluid Mechanics: Realism in fluid statics problems is achieved by utilizing extensive lists of fluid and solid densities.
-This allows for the accurate calculation of hydrostatic pressure and buoyancy forces for a wide range of scenarios, such as a steel object submerged in seawater.
-The parameterization is context-aware, selecting dense fluids like Mercury for manometer problems while using lighter fluids like oil or water for pipe flow scenarios.
-  • Vibrations and Acoustics: For problems involving dynamic systems, parameters such as mass, spring stiffness, and damping coefficients are randomized within ranges that represent plausible physical objects, from small mechanical components to larger structures.
-Additionally, the amplitudes and frequencies of harmonic forcing functions are chosen in relation to the system’s natural frequency to ensure the dynamic response is well-defined and clearly demonstrates key physical phenomena like resonance or beating.
-Appendix F Authoritative Data Sources for Parameterization
-To ensure the physical and engineering realism of EngChain, we built our domain-aware parameterization (discussed in Section 4.4) on authoritative data.
+Appendix E Authoritative Data Sources for Parameterization
+To ensure the physical and engineering realism of EngTrace, we built our domain-aware parameterization on authoritative data.
 We extracted and manually verified extensive lists of physical constants and material properties from the standard handbooks and data sources listed in Table 3.
@@
 Table 3: Primary data sources referenced for engineering parameterization across domains.
-Appendix G LLM Prompt for AI-Assisted Quality Assurance (QA)
-System Prompt
-You are an expert engineering professor and a senior Python developer acting as a peer reviewer for the EngChain benchmark.
+Appendix F Detailed Domain-Aware Parameterization Examples
+To ensure that problems are not just mathematically solvable but are also grounded in physical and engineering realism, EngTrace constrains all generated values by the principles of the domain being tested.
+The following provides specific examples of this methodology for each engineering branch.
+F.0.1 Chemical Engineering
+  • Reaction Kinetics and Stoichiometry: Instead of abstract chemicals, problems select from curated lists of common gas, liquid, and biochemical reactants like Propane, Benzene, and Glucose.
+More importantly, many problems are built around pre-validated, balanced chemical equations to ensure that the fundamental principles of mass conservation and stoichiometry are carefully upheld from the outset.
+  • Thermodynamics: Problems involving thermodynamic calculations use a multi-layered approach to realism.
+For common substances, a complete and consistent set of critical properties ($T_{c},P_{c},V_{c}$) is retrieved.
+For reaction calorimetry, problems utilize tabulated standard heats of formation ($\Delta H_{f}^{\circ}$).
+Crucially, for heat transfer calculations, the model moves beyond constant specific heats by incorporating temperature-dependent heat capacity parameters ($C_{p}/R=A+BT+CT^{2}+DT^{-2}$), ensuring a high degree of physical accuracy across a range of conditions.
+  • Transport Phenomena and Fluid Mechanics: To ensure realism in fluid dynamics problems, templates draw from extensive tables of fluid properties, including the density and viscosity of common liquids and gases.
+The benchmark also includes parameters for non-Newtonian fluids (e.g., Polymer Solutions & Melts etc), allowing it to test more complex, real-world rheological behaviors beyond simple Newtonian assumptions.
+F.0.2 Electrical Engineering
+  • Electromagnetics and Waves: Instead of using an arbitrary number, the phase velocity of a wave is determined by selecting a specific propagation medium from a curated list (e.g., Distilled Water, Polyethylene, Fused Silica etc).
+This ensures that all subsequent calculations for parameters like wavelength or wave number are physically consistent with that material’s properties.
+The generation process is further constrained by fundamental physical constants, including the speed of light in a vacuum ($c_{0}$) and the permittivity of free space ($\epsilon_{0}$).
+  • Signals and Systems: In scenarios bridging continuous and discrete-time domains, such as sampling, the sampling frequency is intentionally set well above the Nyquist rate to model ideal conversion and isolate the effects of the discrete-time system.
+For purely discrete-time problems, parameters are procedurally constrained to test specific concepts.
+For instance, signal frequencies are generated as rational multiples of $\pi$, and downsampling problems are deliberately constructed to either cause or avoid aliasing, directly testing the understanding of frequency folding.
+For fundamental operations like convolution, sequences are kept short and integer-valued to focus the analysis on the computational process.
+  • Digital Communications: For deterministic and random signal analysis, probabilities for discrete random variables are generated as exact rational numbers that sum perfectly to one.
+In digital modulation scheme problems, carrier frequencies are constrained to be integer multiples of the bit rate to simplify analysis, energy per bit ($E_{b}$) and signal-to-noise ratios ($E_{b}/N_{0}$) are generated in physically typical ranges (e.g., picojoules and realistic dB values), and modulation orders ($M$) are restricted to powers of two.
+Furthermore, data rates are assigned practical units such as kbps, Mbps, or Gbps, grounding the problems in contemporary engineering contexts.
+F.0.3 Mechanical Engineering
+  • Mechanics of Materials: Problems are grounded in the behavior of real-world materials by co-selecting properties from comprehensive tables.
+When a material such as “6061-T6 Aluminum” or “Carbon Fiber Reinforced Polymer” is chosen, its physically consistent Young’s Modulus ($E$), Poisson’s Ratio ($\nu$), and Shear Modulus ($G$) are used for all calculations.
+This ensures that problems involving stress, strain, and torsional deformation are realistic.
+Properties are available in both SI (GPa) and US Customary (ksi) units to reflect practical engineering work.
+  • Fluid Mechanics: Realism in fluid statics problems is achieved by utilizing extensive lists of fluid and solid densities.
+This allows for the accurate calculation of hydrostatic pressure and buoyancy forces for a wide range of scenarios, such as a steel object submerged in seawater.
+The parameterization is context-aware, selecting dense fluids like Mercury for manometer problems while using lighter fluids like oil or water for pipe flow scenarios.
+  • Vibrations and Acoustics: For problems involving dynamic systems, parameters such as mass, spring stiffness, and damping coefficients are randomized within ranges that represent plausible physical objects, from small mechanical components to larger structures.
+Additionally, the amplitudes and frequencies of harmonic forcing functions are chosen in relation to the system’s natural frequency to ensure the dynamic response is well-defined and clearly demonstrates key physical phenomena like resonance or beating.
+Appendix G LLM Prompt for Automated AI Pre-Screening
+You are an expert engineering professor and a senior Python developer acting as a peer reviewer for the EngTrace benchmark.
 Your task is to meticulously evaluate a new problem template based on its source code and several example outputs.
@@
 Do not add any explanatory text or markdown formatting around the JSON object.
-User Prompt
 Please evaluate the following engineering problem template.
@@
 ‘‘explanation": "<string, a concise, one-sentence justification for the scores and flag>"
-Appendix H Multi-Axis Rubric for LLM-Based Template Validation
-As part of our AI-Assisted Quality Assurance (QA) workflow, the LLM Peer Reviewer evaluates each template against a set of criteria.
-The multi-axis rubric used for this validation, detailing each scoring category and its description, is defined in Table 5.
+Appendix H Multi-Axis Rubric for Automated AI Pre-Screening
+During automated AI pre-screening, judges in the AI Tribunal evaluate templates against a set of criteria. The evaluation rubric, detailing scoring categories and descriptions, is provided in Table 4.
 Score Category Description
@@
 Explanation A concise, one-sentence justification for the scores and the flag.
-Table 5: Multi-axis rubric used for LLM-based template validation. Numerical criteria are rated on a 1–5 scale, and the Human Review Flag is a binary (true/false) indicator.
-Appendix I LLM-as-Judge Error Categories
-During the second stage of our evaluation framework (Section 5), the LLM-as-a-Judge system performs a qualitative error diagnosis on failed reasoning steps.
-This provides deeper insight than a quantitative score alone by distinguishing between fundamental reasoning failures and valid alternative solutions.
-The five distinct error categories used for this classification are defined in Table 6.
-Error Category Description
-Conceptual Error The model applied the wrong scientific principle or formula (e.g., used addition instead of subtraction).
-Calculation Error The model used the correct formula but made a mathematical mistake (e.g., $2\times 3=5$).
-Input Error The model used the correct formula but pulled the wrong number from the problem context or a previous step.
-Alternative Correct The model’s step is valid and logically sound but follows a different method or phrasing.
-Other The model’s step is nonsensical, irrelevant, a hallucination, or contains only formatting errors.
-Table 6: Error taxonomy used to categorize model mistakes during LLM-as-Judge evaluation.
-Appendix J LLM Prompt for Automated Error Analysis (LLM-as-a-Judge)
+Table 4: Multi-axis rubric used for AI Tribunal based template validation. Numerical criteria are rated on a 1–5 scale, and the Human Review Flag is a binary (true/false) indicator.
+Appendix I Human Certification Interface
+To facilitate the rigorous Stage 2 evaluation, we developed a custom web-based annotation tool using the Streamlit framework.
+This interface, illustrated in Figure 7, was designed to streamline the expert review process while ensuring data consistency for statistical analysis.
+Figure 7: The EngTrace Expert Annotation Interface.
+The dashboard features a dual-view layout for verifying source code against rendered traces, enabling experts to input structured quality scores and definitive admissibility decisions.
+The workflow within the interface proceeds as follows:
+  1. 1.
+    Dual-View Inspection: The main panel displays the rendered natural language Question and step-by-step Solution Trace exactly as they appear in the final dataset.
+Simultaneously, reviewers can toggle a tab to inspect the underlying Python Source Code, allowing them to verify the algorithmic logic alongside the output.
+  2. 2.
+    Dynamic Re-sampling: To verify consistency beyond a single static snapshot, the interface allows reviewers to interactively regenerate instances.
+This enables experts to stress-test the parameterization logic and confirm that physical constraints hold across multiple random seeds.
+  3. 3.
+    Ordinal Quality Scoring: Reviewers evaluate the template across three dimensions (Physical Plausibility, Mathematical Correctness, and Pedagogical Clarity) using interactive sliders (1–5).
+These values provide the ordinal data required for calculating Krippendorff’s Alpha ($\alpha$).
+  4. 4.
+    Binary Certification Gate: The “Certification” toggle forces a definitive Approve or Reject decision, generating the categorical data for Fleiss’ Kappa ($\kappa$).
+  5. 5.
+    Mandatory Feedback Loop: If a template is rejected, the interface enforces a mandatory text input field for specific feedback (e.g., "The mole balance assumption is invalid for variable volume").
+This feedback is logged and routed back to the authors for immediate correction.
+Appendix J Template Validation Results
+J.1 Stage 1: AI Tribunal Impact
+In the initial audit, the AI Tribunal identified 26 Critical Failures and 6 Controversial templates.
+These templates were manually refactored based on the specific failure modes identified by the models (e.g., unit mismatches or unrealistic material properties) until a 100% pass rate was achieved; representative examples of these identified failures and their associated reasoning are provided in Table 5.
+Template Metadata Median Score Votes Model Model Feedback Summary
+adiabatic_flame_temperature Chemical Engineering Thermodynamics 2 2/3 GPT-5 Structure clear but temperatures too low; likely $C_{p}$ polynomial or unit mismatch in energy balance.
+Claude 4.5 Opus Method sound but values physically inaccurate (e.g., Acetylene $\sim$1493K vs $\sim$2500K); suspects data parameter issues.
+Gemini 3 Pro Critical unit error identified in heat capacity integration (incorrect application of gas constant $R$).
+multi_segment_rod Mechanical Engineering Mechanics of Materials 3 2/3 GPT-5 Equations correct but ignores yield limits, resulting in unrealistically high stresses.
+Claude 4.5 Opus Physics generally sound, but questions the validity of applying large tensile loads to concrete.
+Gemini 3 Pro Identifies physically impossible scenarios where brittle materials are subjected to loads exceeding fracture limits.
+Table 5: Qualitative Feedback from the AI Tribunal. Examples of critical failures identified during initial screening, categorized by model-specific technical critiques.
+J.2 Stage 2: Inter-Annotator Agreement
+We measured agreement among domain experts per branch using Fleiss’ Kappa ($\kappa$) for binary admissibility and Krippendorff’s Alpha ($\alpha$) for ordinal quality metrics.
+As shown in Table 6, we observed perfect agreement ($\kappa=1.0$) across all domains regarding the final inclusion of templates.
+This unusually high consensus suggests that the AI Tribunal based stage 1 successfully acted as a rigorous filter, ensuring that only high-quality candidates reached the human review stage.
+Domain Templates Fleiss’ $\kappa$ Interpretation
+Chem. Eng. 30 1.00 Perfect Agreement
+Elec. Eng. 30 1.00 Perfect Agreement
+Mech. Eng. 30 1.00 Perfect Agreement
+Table 6: Binary Admissibility Consensus (Fleiss’ Kappa). Agreement on whether a template should be Included/Rejected.
+Krippendorff’s Alpha ($\alpha$)
+Domain
+Physical
+Plausibility
+Math.
+Correctness
+Pedagogical
+Clarity
+Chem. Eng. 0.224 1.00 0.457
+Elec. Eng. 0.372 1.00 0.332
+Mech. Eng. 0.647 1.00 0.078
+GLOBAL 0.437 1.00 0.269
+Table 7: Ordinal Quality Consistency (Krippendorff’s Alpha). Agreement on 1–5 scoring across three axes.
+For the ordinal quality ratings, Table 7 reveals a clear distinction between objective and subjective metrics.
+Mathematical Correctness achieved perfect agreement ($\alpha=1.0$), reflecting the objective nature of verifying derivations.
+Physical Plausibility showed moderate agreement ($\alpha\approx 0.44$), as experts occasionally debated edge cases in parameter realism.
+Conversely, Pedagogical Clarity exhibited lower agreement ($\alpha\approx 0.27$), highlighting the inherent subjectivity in evaluating educational "style."
+J.3 AI-Human Alignment
+We evaluated the performance of the AI Tribunal (Stage 1) against the Human Experts (Stage 2) using two key metrics:
+  1. 1.
+    Filter Safety (False Positive Rate): The AI Tribunal demonstrated perfect safety (FPR $0.00\%$), as zero templates approved by the AI were subsequently rejected by humans (Target $10\%$ achieved).
+  2. 2.
+    Scoring Alignment (Spearman’s Rho): The calculated Spearman’s Rho was low ($\rho=0.065,p=0.54$), indicating weak rank correlation.
+This is likely an artifact of range restriction: since the AI filter was highly effective, the surviving templates all clustered near the maximum quality score (4–5/5).
+With variance effectively eliminated (a "ceiling effect"), rank correlation becomes statistical noise rather than a meaningful signal of misalignment.
+Appendix K Dataset Statistics
+Structural Balance and Granularity.
+Each branch is represented by exactly 30 templates, ensuring that no single discipline dominates the aggregate performance metrics.
+This uniform top-level distribution controls for branch-specific bias.
+At the granular level, the number of Areas (6–7 per branch) reflects the specific curricular density of each discipline; for instance, Mechanical Engineering consolidates multiple sub-topics into broad areas like Fluid Mechanics, resulting in slightly fewer total areas (6) compared to Chemical and Electrical (7), without sacrificing concept coverage.
+Difficulty Distribution.
+The dataset difficulty is intentionally stratified to mirror curricular progression.
+We prioritize foundational and applied reasoning (Level 1 & 2: $\approx$75%) to test core engineering laws and procedural workflows, while reserving the most complex synthesis tasks for the advanced tier (Level 3: $\approx$25%).
+This weighting ensures models are evaluated primarily on standard practice while still distinguishing elite capabilities in multi-physics integration.
+Table 8 presents the detailed statistical breakdown.
+Branch Structure Difficulty Level Total
+Dom. Areas Easy Int. Adv.
+Chemical 3 7 13 9 8 30
+Electrical 3 7 11 12 7 30
+Mechanical 3 6 10 13 7 30
+Total 9 20 34 34 22 90
+Table 8: Dataset Statistics by Branch.
+The dataset features a perfectly balanced distribution of 30 templates per branch, with difficulty levels scaled to test both foundational knowledge and advanced synthesis.
+Appendix L LLM Prompt for Multi-Model Heuristic Verification
 You are an expert engineering professor acting as an automated evaluator.
@@
 Do not include any introductory text, concluding remarks, or markdown formatting.
-Error Categories: ‘‘Conceptual Error": The model applied the wrong scientific principle or formula (e.g., used addition instead of subtraction).
-‘‘Calculation Error": The model used the correct formula but made a mathematical mistake (e.g., 2 * 3 = 5).
-‘‘Input Error": The model used the correct formula but pulled the wrong number from the problem context or a previous step.
+Error Categories: ‘‘Conceptual Error": The model applied the wrong scientific principle or formula.
+‘‘Calculation Error": The model used the correct formula but made a mathematical mistake.
 ‘‘Alternative Correct": The model’s step is valid and logically sound, but follows a different method or phrasing than the ground-truth step.
@@
 {"error_category": "...", "explanation": "..."}
-Appendix K Evaluated Model Details
-This appendix provides a comprehensive overview of all Large Language Models discussed in our experimental framework (Section 6).
-As outlined in the main text, our evaluation suite is structured around a four-way taxonomy to enable a structured comparison across different training paradigms.
-These four categories are: (1) Frontier Proprietary LLMs, (2) Science / Engineering Enhanced LLMs, (3) Math Enhanced LLMs, and (4) General Purpose Open LLMs.
-Table 7 provides the complete list of models within each category, detailing their developing organization, parameter size (where public), backbone model (for fine-tuned variants), and source identifier (either an API endpoint or a HuggingFace repository path).
+Appendix M Evaluation Framework Validation
+To ensure that the EngTrace evaluation framework reflects expert engineering judgment, we conducted a rigorous human-alignment study.
+We established a validation set of 20 instantiated questions, randomly sampled with stratification across all three branches and difficulty levels, and generated responses from five representative models: deepseek-reasoner, gemini-3-pro-preview, gemma-3-27b-it, qwen2.5-math-7b, and Meta-Llama-3.1-8B-Instruct.
+This procedure resulted in 100 randomized and anonymized responses.
+Three subject matter experts per branch independently scored them on a 1–5 Likert scale.
+Our initial analysis revealed a strong correlation between human-rated reasoning quality and final answer accuracy (Spearman’s $\rho=0.9130$, $p<0.05$), confirming that coherent reasoning traces are highly predictive of numerical correctness.
+Consequently, we prioritized reasoning process quality as the primary dimension for optimizing our automated evaluator.
+M.1 Expert Evaluation Rubrics
+The following rubrics were utilized by subject matter experts to evaluate the 100 validation responses.
+Experts provided independent scores for reasoning and accuracy on a 1–5 Likert scale as detailed in Table 9 and Table 10.
+Score Rating Description
+5 Excellent Flawless derivation; follows all physical laws and engineering principles.
+4 Good Correct reasoning with minor notation inconsistencies or suboptimal paths.
+3 Fair Generally sound logic; minor errors in parameter grounding or arithmetic.
+2 Poor Significant logical gaps or incorrect governing equations for the domain.
+1 Deficient Nonsensical, physically impossible, or complete hallucination of concepts.
+Table 9: Expert rubric used for assessing the quality of the reasoning process.
+Score Rating Description
+5 Excellent Numerically correct within tolerance including correct units.
+4 Good Correct value but missing or incorrectly formatted units.
+3 Fair Off by a small margin due to rounding; correct magnitude.
+2 Poor Numerically incorrect by a significant margin or order of magnitude.
+1 Deficient Entirely wrong or absent.
+Table 10: Expert rubric for assessing final answer accuracy.
+M.2 Ablation Study Results
+We systematically assessed various evaluator configurations to optimize alignment with these expert ratings.
+Table 11 shows the Spearman’s Rank Correlation ($\rho$) between automated pipeline scores and human judgment.
+The results demonstrate that the combination of an AI Tribunal ($\tau=0.8$) and numeric step-matching yields the highest alignment ($\rho=0.632$).
+This alignment significantly degrades upon the removal of the Tribunal ($\rho=0.411$), confirming its necessity for capturing valid alternative derivations.
+Furthermore, the poor performance of standard NLP metrics ($\rho=0.325$) reinforces the need for the domain-specific numerical verification employed in EngTrace.
+Technical Configuration Alignment ($\rho$)
+Tribunal ($\tau=0.8$) + Numeric Step-Matching 0.632
+Tribunal ($\tau=0.95$) + Numeric Step-Matching 0.581
+Tribunal ($\tau=0.6$) + Numeric Step-Matching 0.525
+No AI Tribunal; Strict Ground-Truth Matching 0.411
+Tribunal + Step-Matching (Numbers ignored) 0.469
+ROUGE-L / BERTScore (Standard NLP metrics) 0.325
+Table 11: Ablation results showing the Spearman’s rank correlation ($\rho$) between automated reasoning scores and human expert ratings.
+The data highlights the necessity of the AI Tribunal and numerical verification.
+Appendix N Evaluated Model Details
+Table 12 details the 24 models evaluated in our experiments (Section 5).
+We report the developing organization, parameter size, backbone architecture (where applicable), and the specific access identifier used to ensure reproducibility.
 Model Organization Size Backbone Source
 Frontier Proprietary LLMs
-GPT-5 OpenAI N/A gpt-5-2025-08-07
-GPT-4.1 OpenAI N/A gpt-4.1-2025-04-14
+GPT-5 Base OpenAI N/A gpt-5-base-2025-08-07
 GPT-5 Mini OpenAI N/A gpt-5-mini-2025-08-07
+GPT-4.1 Base OpenAI N/A gpt-4.1-2025-04-14
 GPT-4.1 Mini OpenAI N/A gpt-4.1-mini-2025-04-14
@@
 Claude Sonnet 3.7 Anthropic N/A claude-3-7-sonnet-20250219
-Gemini 2.5 Pro Google DeepMind N/A gemini-2.5-pro-2025
-Gemini 2.5 Flash Google DeepMind N/A gemini-2.5-flash-2025
-DeepSeek V3.2 DeepSeek N/A deepseek-v3.2-20250929
-DeepSeek V3.1 DeepSeek N/A deepseek-v3.1-20250922
-DeepSeek R1 DeepSeek N/A deepseek-r1-20250120
-Grok 4 Heavy xAI N/A grok-4-heavy-0709
-Grok 4 Fast xAI N/A grok-4-fast-reasoning
-Science / Engineering Enhanced LLMs
-Galactica Meta AI 120B facebook/galactica-120b
-Minerva Google DeepMind 540B deepmind/minerva
-ChemLLM Shanghai AI Lab 13B meta-llama/Llama-2-13B ShanghaiAILab/ChemLLM-13B
-SciLit01 TII 7B Qwen/Qwen2-7B-Instruct tiiuae/SciLit-01-7B
-SciAgent OpenBioLLM Team 8B mistralai/Mistral-7B-v0.3 OpenBioLLM/SciAgent-8B
+Gemini 3 Pro Google N/A gemini-3-pro-001
+Gemini 2.5 Pro Google N/A gemini-2.5-pro-002
+Gemini 2.5 Flash Google N/A gemini-2.5-flash-002
+DeepSeek V3 DeepSeek N/A deepseek-v3
+DeepSeek R1 DeepSeek N/A deepseek-reasoner
+General Purpose Open LLMs
+Llama 3.1 70B Meta 70B meta-llama/Llama-3.1-70B-Instruct
+Llama 3.1 8B Meta 8B meta-llama/Llama-3.1-8B-Instruct
+Qwen 2.5 72B Alibaba Cloud 72B Qwen/Qwen2.5-72B-Instruct
+Qwen 2.5 14B Alibaba Cloud 14B Qwen/Qwen2.5-14B-Instruct
+Qwen 2.5 7B Alibaba Cloud 7B Qwen/Qwen2.5-7B-Instruct
+Qwen 3 8B Alibaba Cloud 8B Qwen/Qwen3-8B-Instruct
+Gemma 3 27B Google 27B google/gemma-3-27b-it
+Gemma 2 9B Google 9B google/gemma-2-9b-it
 Math Enhanced LLMs
-WizardMath WizardLM Team 7B mistralai/Mistral-7B-v0.1 WizardLMTeam/WizardMath-7B-V1.1
-MetaMath MetaMath Project 7B EleutherAI/llemma-7b meta-math/MetaMath-7B-V1.0
-Mathstral Mistral AI 7B mistralai/Mistral-7B-v0.1 mistralai/Mathstral-7B-v0.1
-Qwen2.5-Math Qwen Team 7B Qwen/Qwen2.5-7B Qwen/Qwen2.5-Math-7B-Instruct
-General Purpose Open LLMs
-LLaMA 3.1 Meta 8B meta-llama/Llama-3.1-8B
-Qwen 2.5 Alibaba Qwen Team 7B Qwen/Qwen2.5-7B-Instruct
-Qwen 3 Alibaba Qwen Team 8B Qwen/Qwen3-8B
-Table 7: Overview of all models evaluated, grouped by functionality: (1) frontier proprietary LLMs representing current performance upper bounds, (2) science/engineering-enhanced models trained on technical corpora, (3) math-enhanced models for quantitative reasoning, and (4) general-purpose open LLMs providing broad foundations.
+Qwen2.5-Math Qwen Team 7B Qwen2.5-7B Qwen/Qwen2.5-Math-7B-Instruct
+Mathstral Mistral AI 7B Mistral-7B mistralai/Mathstral-7B-v0.1
+WizardMath WizardLM Team 7B Mistral-7B WizardLMTeam/WizardMath-7B-V1.1
+MetaMath MetaMath Project 7B Llemma-7B meta-math/MetaMath-7B-V1.0
+Table 12: Overview of the 24 models evaluated, grouped into (1) frontier proprietary, (2) general-purpose open, and (3) math-enhanced categories.
 For proprietary models, we list the internal or API model identifiers; for open models, we include the HuggingFace source.
-Appendix L AI-Assisted QA Workflow: Template Examples
-This appendix provides visual evidence of the LLM Peer Reviewer’s output for an automatically approved template (provided in Figure 9) and a flagged template (provided in Figure 10), illustrating the function of our automated Quality Assurance (QA) process (Section 4.5).
-Figure 9: LLM Peer Review result for an Approved Template (template_rackett_equation_volume).
-The template, designed for Chemical Engineering Thermodynamics, receives the maximum score of 5/5 across all core categories (Physical Plausibility, Mathematical Correctness, and Pedagogical Clarity), indicating that the problem structure is robust and ready for benchmark inclusion.
-Figure 10: LLM Peer Review result for a Flagged Template (template_euclidean_distance_binary).
-The model detected a fundamental flaw, assigning a score of 3/5 for both Plausibility and Correctness.
-The explicit explanation flags a “calculation error that sets sqrt(Eb) to zero,” which immediately forces the template out of the automated pipeline and into human review for manual correction.
-Appendix M Detailed LLM-as-Judge Error Analysis
-This appendix provides the full, fine-grained breakdown of the qualitative error analysis performed on Gemini 2.5 Pro, using GPT-4o as the judge.
-The following tables detail the distribution of all 4,118 reasoning steps that were flagged by the quantitative verifier as a mismatch from the ground-truth solution.
-Table 8 breaks down the errors by engineering branch,  Table 9 shows the distribution by problem difficulty, and Figure 11 provides a proportional breakdown of error types across all nine engineering domains.
-Branch Alternative Correct Conceptual Error Calculation Error Input Error Other Total
-Chemical 1353 206 67 17 43 1686
-Electrical 827 221 23 24 94 1189
-Mechanical 865 207 63 40 67 1242
-Total 3045 634 153 81 204 4117
-Table 8: Branch-level Error Analysis for Gemini 2.5 Pro.
-This table shows the distribution of 4,117 flagged reasoning mismatches, categorized by the judge (GPT-4o), across the three main engineering branches.
-Difficulty Alternative Correct Conceptual Error Calculation Error Input Error Other Total
-Easy 1088 162 52 15 61 1378
-Intermediate 1178 204 65 51 73 1571
-Advanced 779 268 36 15 70 1168
-Total 3045 634 153 81 204 4117
-Table 9: Difficulty-level Error Analysis for Gemini 2.5 Pro.
-This table shows the distribution of 4,117 flagged reasoning mismatches, categorized by the judge (GPT-4o), across the three problem difficulty levels.
-Figure 11: Domain-level Error Analysis for Gemini 2.5 Pro.
-This 100% stacked bar chart illustrates the proportional distribution of error types across all nine engineering domains.
-This visualization highlights how the ratio of error types changes by domain; for instance, Fluid Mechanics and Signals & Systems show a much larger proportion of Conceptual Errors and Other errors compared to Thermodynamics or Reaction Kinetics, where Alternative Correct solutions dominate.
+Appendix O Inference Prompt and Output Processing
+O.1 Standardized Inference Prompt
+To ensure consistent and parsable outputs across all models, we employed a standardized zero-shot system prompt.
+This prompt explicitly instructs the model to structure its response with specific headings and to format the final result for automated extraction.
+You are an expert engineer. Solve the following
+problem by providing a detailed, structured
+solution. Use the exact headings and
+formatting provided below.
+# Given
+List all known variables and their
+values with units.
+# Find
+State the variable(s) to be calculated.
+# Formulae
+Write down all necessary governing equations
+before substituting any values.
+# Solution
+Provide a step-by-step calculation. Each step
+must start on a new line and be formatted exactly
+as ’**Step X:**’, where X is the step number.
+Show the substitution of values into
+the formulae clearly.
+# Final Answer
+State the final numerical result with its units
+in the format: **Answer:** [value] [units]
+O.2 Output Parsing Logic
+To process the model outputs generated by the prompt above, we employ a heuristic parsing pipeline:
+  1. 1.
+    Step Segmentation: We use regular expressions to identify line-start markers matching either the Step <number> format (e.g., **Step 1**) or the numbered-list format (e.g., 1.).
+A subsequent regex cleans these prefixes to isolate the reasoning text.
+  2. 2.
+    Value Extraction: We extract numerical values from each step and the final answer using a three-priority heuristic:
+    • Priority 1: Matching the explicit **Answer:** tag defined in the prompt.
+    • Priority 2: Identifying the last number following an equals sign (e.g., V = 100.0).
+    • Priority 3: Extracting the last standalone number in the segment.
+  3. 3.
+    Fallback: If no overall **Answer:** tag is found, the final numerical value extracted from the last reasoning step is used for verification.